### v1

# The Euler class of the normal bundle of a Seifert fibration and horizontal foliations

Abstract.
For Seifert fibred manifolds with orientable base orbifolds, we establish a necessary and sufficient condition for the Euler class of the normal bundle of the Seifert fibration to vanish. When the base orbifold is hyperbolic, we also provide a second proof of this condition from the perspective of discrete faithful representations of Fuchsian groups.
As an application, we present infinitely many Seifert fibred rational homology spheres that admit co-oriented taut foliations but none with vanishing Euler class. In the context of the $L$-space conjecture, these provide examples of rational homology spheres that admit co-oriented taut foliations (and hence are not $L$-spaces) and have left-orderable fundamental groups yet none of the left-orders arise directly from the universal circle actions associated to co-oriented taut foliations.
Steven Boyer was partially supported by NSERC grant RGPIN 2024-06196
Ying Hu was partially supported by NSF grant DMS-2409398
2010 Mathematics Subject Classification. Primary 57M25, 57M50, 57M99
Key words: Seifert fibred manifold, normal bundle, Euler class, horizontal foliation
1. Introduction
In this article, we study the triviality of the normal bundle $\nu_{M}$ to the Seifert fibres of a Seifert fibration on a $3$-manifold $M$ with orientable base orbifold. We will refer to $\nu_{M}$ as the normal bundle of the Seifert fibration. As an oriented $2$-plane bundle, $\nu_{M}$ is trivial if and only if its Euler class $e(\nu_{M})\in H^{2}(M)$ vanishes.
We denote a closed oriented Seifert fibred manifold $M$ with orientable base orbifold by $M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, where $b\in\mathbb{Z}$ and $b_{i}\geq 2$ for $i=1,\ldots,n$. See §2.1 for a detailed explanation of the notation. We denote the base orbifold by $\mathcal{B}$; this is a closed orientable surface $|\mathcal{B}|$ of genus $g\geq 0$ with $n\geq 0$ cone points of orders $b_{1},...,b_{n}$. Recall that the Euler characteristic $\chi(\mathcal{B})$ of $\mathcal{B}$ is $\chi(|\mathcal{B}|)-n+\sum_{i=1}^{n}\frac{1}{b_{i}}$, and the Euler number of $M$ is $e(M)=b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}$.
Here is our main result.
Theorem 1.1.
Given a closed oriented Seifert fibred manifold $M=M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, the Euler class of the normal bundle $\nu_{M}$ of the Seifert fibration vanishes if and only if there exists $m\in\mathbb{Z}$ such that
  1. (1)
    $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i}),\text{ for }i=1,\cdots,n,\;\text{and}$
  2. (2)
    $me(M)=\chi(\mathcal{B})$.
Remark 1.2.
It is often easy to identify an initial $m\in\mathbb{Z}$ that satisfies the congruence equations $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i})$ for $i=1,\cdots,n$. Any other integer $m^{\prime}$ satisfying these congruences can be written as $m^{\prime}=m+rl$, where $l$ is the least common multiple of $b_{1},b_{2},\ldots,b_{n}$ and $r\in\mathbb{Z}$. Hence when $n>0$, $e(M)\neq 0$ and $\chi(\mathcal{B})\neq 0$, the condition $m\,e(M)=\chi(\mathcal{B})$ in Theorem 1.1 can be replaced by
$\frac{\chi(\mathcal{B})}{e(M)}\text{ is an integer congruent to }m\;(\operatorname{mod}l)$
Next, we deduce the following easily verifiable necessary condition for $\nu_{M}$ to be a trivial bundle:
Corollary 1.3.
Given a closed oriented Seifert fibred manifold $M$, if the Euler class of the normal bundle $\nu_{M}$ of the Seifert fibration vanishes, then either $M$ is an oriented $S^{1}$-bundle over the torus or one of the following holds:
  1. (1)
    $e(M)=\chi(\mathcal{B})=0$,
  2. (2)
    $e(M)\neq 0,\,\chi(\mathcal{B})\neq 0,$ and
    $\frac{\chi(\mathcal{B})}{e(M)}\in\mathbb{Z}.$
Note that if $M$ is an oriented $S^{1}$-bundle over the torus, then its normal bundle is always trivial (see Remark 2.2).
By abelianizing the presentation (3.0.1) of $\pi_{1}(M)$ obtained by viewing $M$ as doing a Dehn filling of a trivial $S^{1}$-bundle over an $n$-punctured orientable surface of genus $g$ as in §2.1, one can show that if $T_{1}(M)$ denotes the torsion subgroup of $H_{1}(M)$, then
$|T_{1}(M)|=b_{1}\cdots b_{n}|e(M)|,$
where we take the convention that $b_{1}b_{2}\cdots b_{n}=1$ when $n=0$. Thus, it follows from Corollary 1.3 that:
Corollary 1.4.
Consider the Seifert fibred manifold $M=M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ and suppose that the Euler class of the normal bundle $\nu_{M}$ of the Seifert fibration vanishes. Then the integer $b_{1}\ldots b_{n}\chi(\mathcal{B})$ is a multiple of $|T_{1}(M)|$. Hence, if $M$ is a rational homology $3$-sphere, then $b_{1}\ldots b_{n}\chi(\mathcal{B})$ is a multiple of $|H_{1}(M)|$.
Our initial motivation for studying the Euler class of $\nu_{M}$ comes from the study of horizontal foliations on Seifert fibred manifolds. By definition, a horizontal foliation is transverse to the Seifert fibration and as such, their tangent bundles are isomorphic to $\nu_{M}$. Hence, their Euler classes coincide, at least up to sign.
It is predicted by the $L$-space conjecture [3, 13] that if a closed, orientable, prime $3$-manifold admits a co-orientable taut foliation, then its fundamental group is left-orderable. One of the main tools for proving left-orderability of the fundamental group in the presence of a co-orientable taut foliation is to use Thurston’s universal circle action. See [7, 4] for the details. However, this approach requires the Euler class of the taut foliation to vanish. Although it was expected that examples of $3$-manifolds admitting co-orientable taut foliations but none with vanishing Euler class should exist, no published examples were previously known. In this article, we settle the matter by presenting such examples within the family of Seifert fibred rational homology $3$-spheres. Our arguments are based on the following application of Theorem 1.1.
Theorem 1.5.
Suppose that $M=M\bigl(0;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ is a rational homology $3$-sphere which admits a co-orientable taut foliation $\mathcal{F}$. Then $\mathcal{F}$ has zero Euler class if and only if there exists $m\in\mathbb{Z}$ such that
  1. (1)
    $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i}),\text{ for }i=1,\cdots,n,\;\text{and}$
  2. (2)
    $me(M)=\chi(\mathcal{B})$.
The simplest examples of manifolds in which every co-oriented taut foliation has non-zero Euler class arise from Dehn surgery on the trefoil knot. In the following statement, we assume that $p$ and $q$ are nonzero coprime integers.
Theorem 1.6.
The $3$-manifold obtained by $p/q$-surgery along the right-handed trefoil knot admits a co-oriented taut foliation if and only if $p/q<1$. It admits one with zero Euler class if and only if $p/q<1$ and $|q|\equiv 1\pmod{p}$.
We will present more examples of rational homology Seifert fibred manifolds that admit co-oriented taut foliations, but none with vanishing Euler class, in §4.3 (Examples 4.4). In fact, our computations in §4 suggests that most closed Seifert fibred manifolds have nontrivial normal bundle, including the case when the base orbifold is Euclidean (i.e. $\chi(\mathcal{B})=0$). This may seem counterintuitive, since we have pointed out that the normal bundle is always trivial when the base $\mathcal{B}$ is the torus.
We close with an interesting by-product of our analysis.
Theorem 1.7.
Let $M$ be a closed oriented Seifert fibred manifold with orientable base orbifold $\mathcal{B}$. If $\mathcal{B}$ is hyperbolic or Euclidean, and $e(\nu_{M})=0$, then $M$ admits a horizontal foliation.
In contrast, closed Seifert fibred manifolds with spherical base orbifolds have finite fundamental groups and hence cannot admit taut foliations. Yet, in §4.1, we will show examples of such manifolds for which $e(\nu_{M})$ vanishes.
Plan of the paper
In Section 2 we explain our notation for Seifert manifolds and prove Theorem 1.1, which gives necessary and sufficient conditions for the Euler class of the normal bundle of a Seifert fibration to vanish. We also prove Corollary 1.3 in this section. In Section 3 we present a different perspective on Theorem 1.1 and Corollary 1.3 in the case that the base orbifold is hyperbolic through the use of discrete faithful $PSL(2,\mathbb{R})$-representations of $\pi_{1}(\mathcal{B})$. As a consequence of this approach, we prove the hyperbolic case of Theorem 1.7. Finally, in Section 4 we prove Theorem 1.5 and Theorem 1.6 and provide examples of Seifert fibred manifolds whose normal bundles are (non-)trivial. We classify which occurs when $\mathcal{B}$ is spherical in §4.1 and is Euclidean in §4.2. The Euclidean case of Theorem 1.7 is dealt with at the end of §4.2. The case that $\mathcal{B}$ is hyperbolic is analysed in §4.3. Finally, in §4.4, we illustrate how the conditions of Theorem 1.1 can be used to verify the known fact that $e(\nu_{M})=0$ when $|H_{1}(M)|=1,2$.
Acknowledgements
We thank Nathan Dunfield for bringing up the question of finding examples of rational homology spheres that admit co-oriented taut foliations but none with vanishing Euler class, during his lecture at the workshop “Low Dimensional Topology and Floer Theory” in Montreal, August 18-29, 2025, which inspired us to investigate the Euler class of horizontal foliations in Seifert fibred manifolds. The second and third named authors would like to thank the Simons Foundation, the CRM, and CIRGET for their support during the preparation of this article.
2. The Euler class of the normal bundle of a Seifert fibration
Let $M$ be a closed, oriented Seifert fibred manifold whose base orbifold has an orientable underlying space. In this section, we compute the Euler class of the normal bundle of the Seifert fibration on $M$ and determine conditions for it to vanish (Theorem 1.1 and Corollary 1.3).
2.1. Seifert notation
We first recall the Seifert notation that we will use to describe $M$.
Consider $M$ constructed by attaching $n+1$ solid tori $N_{i}$, $i=0,1,\dots,n$, to a trivial $S^{1}$-bundle over an orientable surface $F$ with $|\partial F|=n+1$, where we assume that $N_{0}$ is a tubular neighbourhood of a regular fibre and the $N_{i}$ are tubular neighbourhoods of the singular fibres for $i=1,2,\ldots,n$. Denote this trivial $S^{1}$-bundle by $P\cong F\times S^{1}$. Fix orientations on $F$ and on the fibres so that they intersect positively.
Let
$\partial F=\mu_{0}\sqcup\mu_{1}\sqcup\cdots\sqcup\mu_{n}$
with the induced orientation, and
$\partial P=T_{0}\sqcup T_{1}\sqcup\cdots\sqcup T_{n}$
with $\mu_{i}\subset T_{i}$ for each $i=0,1,\cdots,n$. So on each $T_{i}$, we use the basis given by $\{h_{i},\mu_{i}\}$, where $h_{i}$ is given by a Seifert fibre. By our orientation convention, we have $\mu_{i}\cdot h_{i}=1$.
Let $D_{i}$ be a meridional disk of $N_{i}$, oriented so that it intersects the Seifert fibres positively. Under the attaching map, we have $\partial D_{i}=-\bigl(a_{i}h_{i}+b_{i}\mu_{i}\bigr)$ as oriented curves, for some $a_{i}\in\mathbb{Z}$, $b_{i}\in\mathbb{Z}_{>0}$ and $a_{i}$, $b_{i}$ coprime. In fact, since $N_{0}$ is a tubular neighbourhood of a regular fibre, and the $N_{i}$ are tubular neighbourhoods of singular fibres, we have $b_{0}=1$ and $b_{i}\geq 2$.
Let $g\geq 0$ denote the genus of $F$, $b=\frac{a_{0}}{b_{0}}\in\mathbb{Z}$, and denote
$M=M\Bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\Bigr)$
Let $\mathcal{B}$ denote the base orbifold of $M$, which is a genus $g$ closed orientable surface with $n$ cone points of orders $b_{1},\ldots,b_{n}$ respectively. We use $|\mathcal{B}|$ to denote the underlying space of $\mathcal{B}$.
Remarks 2.1.
  1. (1)
    This description of the Seifert fibred manifold is not unique. More precisely, we have that $M\Bigl(g;\,\tfrac{a_{0}}{b_{0}},\ldots,\tfrac{a_{n}}{b_{n}}\Bigr)$ is homeomorphic to $M\Bigl(g;\,\tfrac{a^{\prime}_{0}}{b_{0}},\ldots,\tfrac{a^{\prime}_{n}}{b_{n}}\Bigr)$ by an orientation-preserving homeomorphism preserving fibres if and only if up to reordering,
    $\frac{a_{i}}{b_{i}}-\frac{a^{\prime}_{i}}{b_{i}}\in\mathbb{Z}\quad\text{and}\quad\sum_{i=0}^{n}\frac{a_{i}}{b_{i}}=\sum_{i=0}^{n}\frac{a^{\prime}_{i}}{b_{i}}.$
    By choosing exactly one of the $N_{i}$, namely $N_{0}$, to be a neighbourhood of a regular fibre in the construction above, we are adopting the usual convention of having only one of the $\tfrac{a_{i}}{b_{i}}$ be an integer.
  2. (2)
    Given $M=M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, the sum
    $e(M):=b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}$
    is an invariant of the given Seifert fibring of $M$ called the Euler number.
Remark 2.2.
When $n=0$, the manifold $M=M(g;b)$ is an $S^{1}$-bundle over a closed, orientable surface $B$ of genus $g$ with Euler number $b$. Moreover, the necessary and sufficient condition in Theorem 1.1 simplifies to the existence of $m\in\mathbb{Z}$ such that
(2.1.1) $mb=\chi(B)$
By assumption, the Euler class of the circle bundle $f:M\to B$ is $bu$, where $u$ is a generator of $H^{2}(B)\cong\mathbb{Z}$. It is easy to see that $\nu_{M}$ is the pull-back by $f$ of the tangent bundle $\tau_{B}$ of $B$. Hence,
(2.1.2) $e(\nu_{M})=f^{*}(e(\tau_{B})),$
Let $\langle h\rangle$ denote the central subgroup of $\pi_{1}(M)$ generated by a circle fibre of $f$ and consider the $5$-term exact sequence in cohomology associated to the central extension $1\to\langle h\rangle\to\pi_{1}(M)\to\pi_{1}(B)\to 1$:
$0\to H^{1}(B)\xrightarrow{}H^{1}(M)\to\mathbb{Z}\xrightarrow{\;\delta\;}H^{2}(B)\xrightarrow{\;f^{*}\;}H^{2}(M)$
It follows from [8, Theorem 4] that the image of 1-1 under the transgression homomorphism $\delta$ is the Euler class $bu$ of the circle bundle. Hence $\mbox{kernel}(f^{*})$ is generated by $bu$. Then by (2.1.2), $e(\nu_{M})=0$ if and only $e(\tau_{B})=\pm\chi(B)u$ is a multiple of $b$, which is condition (2.1.1).
We consider the following two special cases:
  1. (1)
    If $M=M(g;b)$ is a rational homology sphere, then the base $B$ must be the $2$-sphere, and $M=M(0;b)$ is a lens space $L(|b|,1)$. In this case, the Euler class $e(\nu_{M})$ vanishes if and only if $b\in\{\pm 1,\pm 2\}$.
  2. (2)
    When $B$ is the torus, (2.1.2) implies that the Euler class $e(\nu_{M})$ is always trivial, since the tangent bundle $T\mathcal{B}$ is trivial,
2.2. Proof of Corollary 1.3
We prove Corollary 1.3 using Theorem 1.1 here and defer the proof of Theorem 1.1 to the next subsection.
Let $\nu_{M}$ denote the normal plane bundle over $M$ to the Seifert fibration, oriented with the induced orientation from those on $M$ and on the fibres.
Proof of Corollary 1.3.
Assume that the Euler class of the normal bundle vanishes.
We first consider the case $n=0$, in which $\mathcal{B}$ is a genus $g$ surface and $M$ is an $S^{1}$-bundle over $\mathcal{B}$ of Euler number $b$. The necessary and sufficient condition in Theorem 1.1 is equivalent to that the existence of $m\in\mathbb{Z}$ such that $mb=\chi(\mathcal{B})$.
If $m=0$, then $\chi(\mathcal{B})=mb=0$ and therefore $M$ is an oriented $S^{1}$-bundle over the torus. On the other hand, if $M$ is not an oriented $S^{1}$-bundle over the torus, then $\chi(\mathcal{B})\neq 0$. Then for the equality $mb=\chi(\mathcal{B})$ to hold, we must have $e(M)=b\neq 0$. Hence
$m=\frac{\chi(\mathcal{B})}{e(M)}\in\mathbb{Z}.$
Next we consider the case $n>0$. Again by Theorem 1.1, there exists $m\in\mathbb{Z}$ such that $ma_{i}\equiv 1\pmod{b_{i}}$ and $me(M)=\chi(\mathcal{B})$. Since $m$ cannot be zero in this case, either (1) or (2) holds. ∎
2.3. Proof of Theorem 1.1
Here we provide a proof of Theorem 1.1. When the base orbifold is hyperbolic, i.e., $\chi(\mathcal{B})<0$, we will give another perspective on Theorem 1.1 and Corollary 1.3 in §3. There we will show that the Euler class of the normal bundle is precisely the pullback of the Euler class of a discrete faithful representation of $\pi_{1}(\mathcal{B})$ in $PSL(2,\mathbb{R})$ under the homomorphism $\pi_{1}(M)\to\pi_{1}(\mathcal{B})$ induced by the natural projection $M\rightarrow\mathcal{B}$  [^1] Since $M$ is irreducible when $\mathcal{B}$ is hyperbolic, we have $H^{2}(M)\equiv H^{2}(\pi_{1}(M))$..
Proof of Theorem 1.1.
To simplify the notation in the proof, we will use $\tfrac{a_{0}}{b_{0}}$ in place of $b$, as explained in §2.1. Thus
$M=M\bigl(0;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)=M\bigl(0;\,\tfrac{a_{0}}{b_{0}},\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$
We will compute the Euler class of the normal bundle, denoted by $e(\nu_{M})\in H^{2}(M)$, using the following exact sequence of the pair $(M,\sqcup_{i}T_{i})$:
$\cdots\to\oplus_{i}H^{1}(T_{i})\xrightarrow{\delta}H^{2}(M,\sqcup_{i}T_{i})\xrightarrow{i^{*}}H^{2}(M)\to 0$
Note that
$H^{2}(M,\sqcup_{i}T_{i})\cong H^{2}(P,\sqcup_{i}T_{i})\oplus H^{2}(N_{1},T_{1})\oplus\cdots\oplus H^{2}(N_{n},T_{n}),$
and by the Künneth formula,
$H^{2}(P,\sqcup_{i}T_{i})\cong H^{1}(F,\partial F)\otimes H^{1}(S^{1})\oplus H^{2}(F,\partial F)$
Recall that $\partial F=\sqcup_{i}\mu_{i}$ with $\mu_{i}\subset T_{i}$ for each $i=0,1,\ldots,n$. Suppose that oriented simple closed curves $\alpha_{r},\beta_{r}$, $r=1,\ldots,g$, together with oriented arcs $c_{j}$ on $F$ that connect the boundary component $\mu_{j}$ to the boundary component $\mu_{j+1}$, $j=0,1,\ldots,n-1$ when $n>0$, represent a set of generators of $H_{1}(F,\partial F)$. Then $H_{2}(P,\sqcup_{i}T_{i})$ is freely generated by the classes of vertical tori $\alpha_{r}\times S^{1}$, $\beta_{r}\times S^{1}$, the horizontal surface $F$ and vertical annuli $c_{j}\times S^{1}$ when $n>0$. We denote their duals in $H^{2}(P,\partial P)$ by
$\alpha_{r}^{*}\otimes S^{*},\quad\beta_{r}^{*}\otimes S^{*},\quad c_{j}^{*}\otimes S^{*},\quad F^{\ast}$
respectively, where $r=1,\ldots,g$ and $j=0,1,\ldots,n-1$.
Finally, let $D_{i}^{*}\in H^{2}(N_{i},T_{i})$ denote the dual of $D_{i},\partial D_{i}]\in H_{2}(N_{i},T_{i})$ for each $i=0,1,\ldots,n$.
In summary, a cohomology class in
$H^{2}(M,\sqcup T_{i})\cong H^{1}(F,\partial F)\otimes H^{1}(S^{1})\oplus H^{2}(F,\partial F)\oplus H^{2}(N_{1},T_{1})\oplus\cdots\oplus H^{2}(N_{n},T_{n})$
can be uniquely expressed as a linear combination of
$\alpha_{r}^{*}\otimes S^{*},\quad\beta_{r}^{*}\otimes S^{*},\quad F^{*},\quad D_{i}^{*}\quad\text{and}\quad c_{j}^{*}\otimes S^{*}\text{ when }n>0$
where $r=1,\ldots,g$, $j=0,1,\ldots,n-1$ and $i=0,1,\ldots,n$.
The proof proceeds in two steps (1) and (2).
  1. (1)
    Find the preimage of the Euler class $e(\nu_{M})$ of the normal bundle $\nu_{M}$ in $H^{2}(M,\sqcup T_{i})$.
    Let $\sigma$ be a nowhere vanishing vector field along $\partial P=\sqcup T_{i}$ that points out of $P$. Since $T_{i}$ is tangent to the Seifert fibres, $\sigma$ is a section of $\nu_{M}$ along $\sqcup T_{i}$. We give a cellular structure on $M$ for which we assume that the vertical surfaces $\alpha_{r}\times S^{1}$, $\beta_{r}\times S^{1}$, $c_{j}\times S^{1}$, and the horizontal surfaces $F$ and $D_{i}$, as well as the $T_{i}$, all belong to the 2-skeleton. We also assume that the curves $\alpha_{r}$, $\beta_{r}$ and the arcs $c_{j}$ on $F$ are in the 1-skeleton.
    From $\sigma$ one can define an obstruction $2$-cochain, denoted by $c_{\sigma}$, that represents the relative Euler class of the normal bundle $\nu_{M}$ in $H^{2}(M,\sqcup T_{i})$. By construction, the image of $c_{\sigma}]$ under $H^{2}(M,\sqcup T_{i})\to H^{2}(M)$ is the Euler class $e(\nu_{M})$. (See [9, Section 2] for instance.)
    To write $c_{\sigma}]$ as a linear combination of $\alpha_{r}^{*}\otimes S^{*}$, $\beta_{r}^{*}\otimes S^{*}$, $c_{j}^{*}\otimes S^{*}$, $F^{*}$, and $D_{i}^{*}$, it suffices to compute its evaluation on the dual homology classes.
    1. (a)
      $c_{\sigma}]$ evaluates to zero on all vertical surfaces. This is because, starting from $\sigma|_{\partial F}$, where $\partial F\subset\partial P=\sqcup T_{i}$, there is no obstruction to extending $\sigma$ to the 1-skeleton of $F$. Since $P\cong F\times S^{1}$ is a trivial $S^{1}$-bundle, one can extend the section further to the vertical surfaces $\alpha_{r}\times S^{1}$, $\beta_{r}\times S^{1}$ and $c_{j}\times S^{1}$ by staying constant along the fibre direction.
    2. (b)
      By the Poincaré-Hopf Theorem, $c_{\sigma}]$ evaluated on a horizontal surface equals its Euler characteristic, i.e.,
      $c_{\sigma}]([D_{i},\partial D_{i}])=1,\quad i=0,1,\ldots,n,$
      and
      $c_{\sigma}]([F,\partial F])=\chi(F)=\chi(|\mathcal{B}|)-(n+1).$
  2. (2)
    Compute the image of $\delta:\oplus_{i}H^{1}(T_{i})\rightarrow H^{2}(M,\sqcup_{i}T_{i})$.
    Let $h_{i}^{*},\mu_{i}^{*}\in H^{1}(T_{i})$ denote the (Hom) dual classes of $h_{i}],[\mu_{i}]\in H_{1}(T_{i})$ for $i=0,1,\ldots,n$. We compute $\delta h_{i}^{*}$ and $\delta\mu_{i}^{*}$, which generate the image of $\delta$.
    1. (a)
      It is easy to see that, since $\delta h_{i}^{*}=h_{i}^{*}\circ\partial$ and $\delta\mu_{i}^{*}=\mu_{i}^{*}\circ\partial$, they vanish on all vertical tori, $\alpha_{r}\times S^{1}$ and $\beta_{r}\times S^{1}$.
    2. (b)
      As $\partial([c_{j}\times S^{1}])=[h_{j+1}]-[h_{j}]$ in $H_{1}(\sqcup_{i}T_{i})$ for $j=0,1,\ldots,n-1$, we have
      $\delta\mu_{i}^{*}([c_{j}\times S^{1}])=0\quad\text{for all }i=0,1,\ldots,n,\text{ and }\ j=0,1,\ldots,n-1,$
      and
      $\delta h_{i}^{*}([c_{j}\times S^{1}])=\begin{cases}1&\text{if }j=i-1\\[6.0pt] -1&\text{if }j=i\\[6.0pt] 0&\text{otherwise}\end{cases}$
    3. (c)
      Since $\partial F=\sqcup_{i}\mu_{i}$, we have $\delta h_{i}^{*}([F,\partial F])=0$ and $\delta\mu_{i}^{*}([F,\partial F])=1$ for all $i=0,1,\ldots,n.$
    4. (d)
      Finally, to compute the values of $\delta h_{i}^{*}$ and $\delta\mu_{i}^{*}$ over $D_{i},\partial D_{i}]\in H_{2}(N_{i},T_{i})$, we note that $\partial[D_{i}]=-(a_{i}[h_{i}]+b_{i}[\mu_{i}])$ by our orientation convention in order to keep $b_{i}>0$ (see §2.1). Then it follows that for $i=0,1,\ldots,n$
      $\delta h_{i}^{*}([D_{i},\partial D_{i}])=h_{i}^{*}(-a_{i}[h_{i}]-b_{i}[\mu_{i}])=-a_{i}$
      and
      $\delta\mu_{i}^{*}([D_{i},\partial D_{i}])=\mu_{i}^{*}(-a_{i}[h_{i}]-b_{i}[\mu_{i}])=-b_{i}.$
In summary, we have
$c_{\sigma}]=\chi(F)F^{*}+\sum_{i=0}^{n}D_{i}^{*}$
The image of $\delta$ is generated by
$\delta\mu_{i}^{*}=F^{*}-\sum_{i=0}^{n}b_{i}D_{i}^{*},\quad\text{ and }\quad\delta h_{i}^{*}=\begin{cases}-c_{0}^{*}\otimes S^{*}-a_{0}D_{0}^{*},&\text{for }i=0,\\[6.0pt] c_{i-1}^{*}\otimes S^{*}-c_{i}^{*}\otimes S^{*}-a_{i}D_{i}^{*},&\text{for }0<i<n,\\[6.0pt] c_{n-1}^{*}\otimes S^{*}-a_{n}D_{n}^{*},&\text{for }i=n.\end{cases}$
if $n>0$; and when $n=0$, the image of $\delta$ is generated by
$\delta\mu_{0}^{*}=F^{*}-b_{0}D_{0}^{*}=F^{*}-D_{0}^{*}\quad\text{and}\quad\delta h_{0}^{*}=-a_{0}D_{0}^{*}=-bD_{0}^{*}.$
The Euler class $e(\nu_{M})=i^{*}([c_{\sigma}])\in H^{2}(M)$ vanishes if and only if $c_{\sigma}]\in\mathrm{Im}(\delta)$. That is, there exist integers $m_{i},n_{i}$, for $i=0,1,\ldots,n$, such that
$\delta\Bigl(\sum_{i}-m_{i}h_{i}^{*}+\sum_{i}n_{i}\mu_{i}^{*}\Bigr)=[c_{\sigma}].$
  1. (1)
    When $n>0$, for the coefficients of $c_{j}^{*}\otimes S^{*}$ to be zero, the $m_{i}$ must be identical. Hence it follows that $c_{\sigma}]\in\mathrm{Im}(\delta)$ if and only if there exist $m,n_{i}\in\mathbb{Z}$ for $i=0,\ldots,n$ such that
    $\begin{cases}ma_{i}-n_{i}b_{i}=1,\mbox{ for }i=0,1,\ldots,n\\[6.0pt] \sum_{i=0}^{n}n_{i}=\chi(F)=\chi(|\mathcal{B}|)-(n+1).\end{cases}$
    Recall that $\tfrac{a_{0}}{b_{0}}=b$. Hence the equation $ma_{0}-n_{0}b_{0}=1$ simplifies to $mb-n_{0}=1$, and therefore, the condition above is equivalent to the existence of $m,n_{i}\in\mathbb{Z}$ for $i=1,\ldots,n$ such that
    $\begin{cases}ma_{i}-n_{i}b_{i}=1,\text{ for }i=1,\ldots,n\\[6.0pt] mb-1+\displaystyle\sum_{i=1}^{n}n_{i}=\chi(F)=\chi(|\mathcal{B}|)-(n+1)\end{cases}$
    For $i\geq 1$, we have
    $n_{i}=\frac{ma_{i}-1}{b_{i}}=\frac{ma_{i}}{b_{i}}-\frac{1}{b_{i}}$
    Hence, the second condition is equivalent to
    $m(b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}})=\chi(|\mathcal{B}|)-n+\sum_{i=1}^{n}\frac{1}{b_{i}}$
    Note that
    $e(M)=b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}\quad\text{and}\quad\chi(\mathcal{B})=\chi(|\mathcal{B}|)-n+\sum_{i=1}^{n}\frac{1}{b_{i}}$
    This completes the proof of the theorem when $n>0$.
  2. (2)
    If $n=0$ (and hence there is no singular fibre), the condition above reduces to the existence of $m\in\mathbb{Z}$ such that
    $mb=\chi(\mathcal{B})$
    From the discussion above, one can easily verify that this is a necessary and sufficient condition for the Euler class to vanish in this special case.
This proves the theorem. ∎
3. Discrete representations of Fuchsian groups and the normal bundle of a Seifert fibration
In this section we provide alternative proofs of Theorem 1.1 and Corollary 1.3 from the perspective of discrete faithful representations of Fuchsian groups. We normalise our notation for oriented Seifert manifolds $M=M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ by further requiring that $\tfrac{a_{i}}{b_{i}}\in(0,1)$ for each $i=1,\ldots,n$. See Remark 2.1.
Figure 1. The generators $\bar{x}_{i}$ in the presentation of $\Gamma=\pi_{1}(\mathcal{B})$
The fundamental group of $M$ has a presentation
(3.0.1) $\hskip 28.45274pt\langle x_{1},x_{2},\ldots,x_{n},w_{1},w_{2},\ldots,w_{g},z_{1},z_{2},\ldots,z_{g},h\;|\;h\mbox{ central},x_{i}^{b_{i}}=h^{-a_{i}}\mbox{ for each }i,$
$\hskip 56.9055pt[w_{1},z_{1}][w_{2},z_{2}]\cdots[w_{g},z_{g}]x_{1}x_{2}\cdots x_{n}=h^{b}\rangle,$
and $\Gamma=\pi_{1}(\mathcal{B})$ has a presentation
(3.0.2) $\hskip 28.45274pt\langle\bar{x}_{1},\ldots,\bar{x}_{n},\bar{w}_{1},\ldots,\bar{w}_{g},\bar{z}_{1},\ldots,\bar{z}_{g}\;|\;\bar{x}_{i}^{b_{i}}=1,[\bar{w}_{1},\bar{z}_{1}]\cdots[\bar{w}_{g},\bar{z}_{g}]\bar{x}_{1}\cdots\bar{x}_{n}=1\rangle$
These two presentations are related via a well-known central extension:
(3.0.3) $E_{M}:1\to\mathbb{Z}\to\pi_{1}(M)\xrightarrow{\;\alpha\;}\Gamma\to 1,$
Here $h$ corresponds to the image in $\pi_{1}(M)$ of $1\in\mathbb{Z}$, $\alpha(x_{i})=\bar{x}_{i}$ for each $i$, $\alpha(w_{j})=\bar{w}_{j}$ for each $j$ and $\alpha(z_{k})=\bar{z}_{k}$ for each $k$.
3.1. The second cohomology group of $\Gamma$
Jankins and Neuman determined a canonical isomorphism between $H^{2}(\Gamma)$ and the abelian group having generators $A_{0},A_{1},\ldots,A_{n}$ and relations $A_{0}=b_{i}A_{i}$ for $\leq i\leq n [11$]. (The case that $g=0$ is discussed in [11] and the general case in [10].) Each element of $H^{2}(\Gamma)$ can be uniquely expressed in the form
$dA_{0}+c_{1}A_{1}+\cdots+c_{n}A_{n},$
where $d,c_{1},\ldots,c_{n}\in\mathbb{Z}$ and $1\leq c_{i}\leq b_{i}-1$ for each value of $i$.
The Euler class $e(E_{M})\in H^{2}(\Gamma)$ is the cohomology class which classifies the extension (3.0.3). In [11] it was shown to be
(3.1.1) $e(E_{M})=-(bA_{0}+a_{1}A_{1}+a_{2}A_{2}+\cdots+a_{n}A_{n})$
This class arises as the image of 1-1 under the transgression homomorphism $\delta$ from the $5$-term exact sequence in cohomology of $E_{M}$:
(3.1.2) $0\to H^{1}(\Gamma)\xrightarrow{}H^{1}(M)\to\mathbb{Z}\xrightarrow{\;\delta\;}H^{2}(\Gamma)\xrightarrow{\;\alpha^{*}\;}H^{2}(M)$
See [8, Theorem 4]. Thus,
(3.1.3) $\mbox{kernel}(\alpha^{*})=\langle e(E_{M})\rangle$
3.2. The normal bundle of the Seifert fibring of $M$ and discrete representations
Since $\mathcal{B}$ is hyperbolic, its universal cover can be identified with $\mathbb{H}^{2}$ and the associated group of deck transformations with a discrete subgroup of $\mbox{Isom}_{+}(\mathbb{H}^{2})=PSL_{2}(\mathbb{R})$. This yields a discrete faithful representation $\bar{\varphi}:\Gamma\to PSL_{2}(\mathbb{R})$.
There is a Seifert fibring $\widehat{M}\to\mathbb{H}^{2}$ lifting $M\to\mathcal{B}$ for which the following diagram commutes
$\widehat{M}=\mathbb{H}^{2}\times S^{1}$$\mathbb{H}^{2}$$M\mathcal{B}$
where $\widehat{M}$ as the total space of an $S^{1}$-bundle over the contractible $\mathbb{H}^{2}$ can be identified with $\mathbb{H}^{2}\times S^{1}$. Hence, we can identify the universal cover of $M$ with $\mathbb{H}^{2}\times\mathbb{R}$, where the vertical lines $\{x\}\times\mathbb{R}$ universally cover the Seifert fibres of $M$. Denote the normal bundle to these vertical lines by $\tilde{\nu}$.
The fundamental group of $M$ acts on $\mathbb{H}^{2}$ via the composition $\varphi=\bar{\varphi}\circ\alpha$, which gives rise to a diagonal action $\psi$ of $\pi_{1}(M)$ on $\widetilde{M}\times\mathbb{H}^{2}$:
$\psi(\gamma)(x,y)=(\gamma\cdot x,\varphi(\gamma)(y))$
This action is equivariant with respect to the projection $\widetilde{M}\times\mathbb{H}^{2}\to\widetilde{M}$, so taking quotients determines a locally-trivial $\mathbb{H}^{2}$-bundle
$M\times_{\psi}\mathbb{H}^{2}=(\widetilde{M}\times\mathbb{H}^{2})/\pi_{1}(M)\to M$
Fix a Riemannian metric on $M$ and pull it back to $\widetilde{M}$. Since the group of deck transformations of the cover $\widetilde{M}\to M$ acts as isometries of this metric, the associated exponential map $\mbox{exp}:T\widetilde{M}\to\widetilde{M}$ is $\pi_{1}(M)$-equivariant, as is the composition
$E(\tilde{\nu}_{M})\xrightarrow{\;\mbox{{\footnotesize exp}}\;}\widetilde{M}=\mathbb{H}^{2}\times\mathbb{R}\xrightarrow{\;\mbox{{\footnotesize projection}}\;}\mathbb{H}^{2}$
Now proceed as in the proof of [2, Lemma A.3], starting from its third paragraph, to show that the $\mathbb{H}^{2}$-bundle $M\times_{\psi}\mathbb{H}^{2}\to M$ is topologically equivalent to the normal bundle $\nu_{M}$ and conclude, as in [2, Proposition A4], that $e(\nu_{M})=e(\varphi)=e(\bar{\varphi}\circ\alpha)$. Thus
(3.2.1) $e(\nu_{M})=\alpha^{*}(e(\bar{\varphi}))$
It is well-known that $e(\bar{\varphi})$, respectively $e(\varphi)$, is the obstruction to lifting the representation $\bar{\varphi}:\Gamma\to PSL_{2}(\mathbb{R})$, respectively $\varphi:\pi_{1}(M)\to PSL_{2}(\mathbb{R})$, to $\widetilde{PSL}_{2}(\mathbb{R})$ [14, Lemma 2].
An expression for $e(\bar{\varphi})$ in terms of the Jankins-Neumann generators $A_{0},A_{1},\ldots,A_{n}$ of $H^{2}(\Gamma)$ can be determined following the method of §3 of [11]. To explain it, recall the central extension $1\to\mathbb{Z}\to\widetilde{PSL}_{2}(\mathbb{R})\to PSL_{2}(\mathbb{R})\to 1$. Let $h\in\widetilde{PSL}_{2}(\mathbb{R})$ correspond to $1\in\mathbb{Z}$.
By construction, the generator $\bar{x}_{i}$ of the presentation (3.0.2) of $\Gamma$ is mapped to a conjugate of a rotation of angle $\frac{2\pi}{b_{i}}$ in $PSL_{2}(\mathbb{R})\leq\mbox{Homeo}_{+}(\mathbb{S}^{1})$ by $\bar{\varphi}$, so lifts to a conjugate of translation by $\frac{1}{b_{i}}$ in $\widetilde{PSL}_{2}(\mathbb{R})\leq\mbox{Homeo}_{+}(\mathbb{R})$. Hence, if
$1\to\mathbb{Z}\to\widetilde{\Gamma}\to\Gamma\to 1$
is the pull-back central extension determined by $\bar{\varphi}$, then $\widetilde{\Gamma}$ has a presentation
$\langle\tilde{x}_{1},\ldots,\tilde{x}_{n},\tilde{w}_{1},\ldots,\tilde{w}_{g},\tilde{z}_{1},\ldots,\tilde{z}_{g}\;|\;\tilde{x}_{i}^{b_{i}}=h,[\tilde{w}_{1},\tilde{z}_{1}]\cdots[w_{g},\tilde{z}_{g}]\tilde{x}_{1}\cdots\tilde{x}_{n}=h^{-b}\rangle$
It is a classic result in the theory of Fuchsian groups that under our assumptions, $b=\chi(|\mathcal{B}|)-n$. As such, the discussion in §3 of [11] implies that
(3.2.2) $e(\bar{\varphi})=(n-\chi(|\mathcal{B}|))A_{0}-A_{1}-A_{2}-\cdots-A_{n}$
Proof of Theorem 1.1 when $\mathcal{B}$ is hyperbolic.
As $\nu_{M}$ is an oriented plane bundle, it is trivial if and only if $e(\nu_{M})=0$. Combining (3.2.1), (3.1.3), (3.1.1), and (3.2.2) shows that $e(\nu_{M})=0$ if and only if there is an integer $m$ such that
(3.2.3) $(\chi(|\mathcal{B}|)-n)A_{0}+A_{1}+\cdots+A_{n}=m(bA_{0}+a_{1}A_{1}+a_{2}A_{2}+\cdots+a_{n}A_{n})$
Consideration of the normal form for elements of $H^{2}(\Gamma)$ described in §3.1 shows that the existence of such an $m$ is equivalent to the existence of an integer $m$ for which the following two conditions hold
  1. (1)
    $ma_{i}\equiv 1\mbox{ (mod $b_{i}$) for }1\leq i\leq n$;
  2. (2)
    $mb+\sum_{i=1}^{n}\left\lfloor\frac{ma_{i}}{b_{i}}\right\rfloor=\chi(|\mathcal{B}|)-n$.
For each $i\in\{1,2,\ldots,n\}$, it follows from the congruence equation $ma_{i}\equiv 1\pmod{b_{i}}$ that
$\left\lfloor\frac{ma_{i}}{b_{i}}\right\rfloor=\frac{ma_{i}}{b_{i}}-\frac{1}{b_{i}}$
Hence, the second equation becomes
$m\left(b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}\right)=\chi(|\mathcal{B}|)-n+\sum_{i=1}^{n}\frac{1}{b_{i}}$
which again simplifies to
$me(M)=\chi(\mathcal{B})$
This completes the proof. ∎
Though we proved Corollary 1.3 as a consequence of Theorem 1.1 in §2.2, we provide a proof of the hyperbolic case of the corollary next which is independent of that theorem.
Proof of Corollary 1.3 when $\mathcal{B}$ is hyperbolic.
Let $l$ denote the least common multiple of the integers $b_{1},b_{2},\ldots,b_{n}$ and note that as $\gcd(\frac{l}{b_{1}},\frac{l}{b_{2}},\ldots,\frac{l}{b_{n}})=1$, we have an epimorphism
$\theta:H^{2}(\Gamma)\to\mathbb{Z},\;A_{0}\mapsto l,A_{i}\mapsto\frac{l}{b_{i}}\mbox{ for }1\leq i\leq n$
Then by (3.1.1),
(3.2.4) $\theta(e(E_{M}))=-l(b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}})=-le(M),$
where $e(M)$ is the Euler number of $M$ (see §2.1).
Next note that (3.2.2) implies that
(3.2.5) $\theta(e(\bar{\varphi}))=l(n-\chi(|\mathcal{B}|)-\sum_{i}\frac{1}{b_{i}})=-l\chi(\mathcal{B})$
Suppose that $e(\nu_{M})=0$. Then (3.2.1) and (3.1.3) imply that there is an integer $m$ such that $e(\bar{\varphi})=me(E_{M})\in H^{2}(\Gamma)$ and therefore $\theta(e(\bar{\varphi}))=m\theta(e(E_{M}))$. Since $\mathcal{B}$ is hyperbolic, $\chi(\mathcal{B})\neq 0$, and therefore (3.2.4) and (3.2.5) imply that $m\neq 0$ and $\chi(\mathcal{B})=me(M)$, which is what was to be proved. ∎
Proof of Theorem 1.7 when $\mathcal{B}$ is hyperbolic.
We saw in (3.2.1) that that $e(\nu_{M})=\alpha^{*}(e(\bar{\varphi}))=e(\bar{\varphi}\circ\alpha)=e(\varphi)$, where $\varphi=\bar{\varphi}\circ\alpha:\pi_{1}(M)\to PSL_{2}(\mathbb{R})$. Hence, if $e(\nu_{M})=0$, then $\varphi$ lifts to a homomorphism $\pi_{1}(M)\to\widetilde{PSL}_{2}(\mathbb{R})\leq\mbox{Homeo}_{+}(\mathbb{R})$ [14, Lemma 2]. It follows that $\pi_{1}(M)$ is left-orderable [6, Theorem 1.1] and therefore, $M$ admits a horizontal foliation [6, Theorem 1.3]. ∎
4. Examples and applications
Recall that if $M$ is a Seifert fibred rational homology $3$-sphere, then the underlying surface of the base orbifold is either $S^{2}$ or $P^{2}$. However, in the latter case, $M$ does not admit a co-orientable taut foliation. (See [6] for instance.)
In this section we use Theorem 1.1 and Corollary 1.3 to analyse when an oriented Seifert manifold $M=M(0;b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$ has a trivial normal bundle and apply this to show that generically, if $M$ has a co-oriented taut foliation, it has none with zero Euler class.
We assume that $n\geq 3$. Then the base orbifold $\mathcal{B}$ of $M$ is spherical if $\chi(\mathcal{B})>0$, Euclidean if $\chi(\mathcal{B})=0$, and hyperbolic if $\chi(\mathcal{B})<0$. Orient $\mathcal{B}$ and endow $\nu_{M}$, the normal bundle to the Seifert fibres of $M$, with the induced orientation. The following lemma is key to our applications.
Lemma 4.1.
Suppose that $M$ is a Seifert fibred rational homology $3$-sphere. Then
$(1)$ Any co-orientable taut foliation on $M$ is horizontal. That is, it can be isotoped to be everywhere transverse to the Seifert fibres of $M$.
$(2)$ The tangent space to any co-orientable taut foliation on $M$ is isomorphic to $\nu_{M}$ as an unoriented bundle. In particular, the Euler class of any co-oriented taut foliation on $M$ agrees, up to sign, with that of $\nu_{M}$.
Proof.
That any co-oriented taut foliation on $M$ must be horizontal is [5, Corollary 7] in the case $n=3$ and [1, Proposition 6.1] in general. Thus (1) holds. Consequently, the tangent bundle of any such foliation is isomorphic to $\nu_{M}$ (as an unoriented bundle), which implies (2). ∎
Theorem 1.5 follows immediately from this lemma and Theorem 1.1.
Example 4.2 (Dehn surgeries on the trefoil knot).
Let $K=T(2,3)$ be the right-handed trefoil knot, and let $\mu$ and $\lambda$ denote the standard meridian and longitude of $K$. There is a natural Seifert fibre structure on the exterior $X=S^{3}\setminus\nu(K)$, whose regular fibre has slope $6\mu+\lambda$ on $\partial X$. Hence, $p/q$-surgery on $K$ yields a small Seifert fibred space whenever $p/q\neq 6$.
It is known that $X(p/q)$ admits co-oriented taut foliations when $p/q<1$ [16, §3]. On the other hand, [9, Theorem 1.7] shows that the Euler class of the co-oriented taut foliation constructed by Roberts in [16] vanishes if and only if
$|q|\equiv 1\pmod{p}$.
Combining Example 4.2 with Lemma 4.1 yields Theorem 1.6.
The following lemma is an immediate consequence of Theorem 1.1.
Lemma 4.3.
If $e(\nu_{M})=0$, then $d=\gcd(b_{i},b_{j})\implies a_{i}\equiv a_{j}\mbox{ {\rm(mod $d$)}}$.
Two sets of Seifert invariants $(b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$, $(b^{\prime},\frac{a_{1}^{\prime}}{b_{1}},\ldots,\frac{a_{n}^{\prime}}{b_{n}})$ are said to be equivalent, written
$(b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})\sim(b^{\prime},\frac{a_{1}^{\prime}}{b_{1}},\ldots,\frac{a_{n}^{\prime}}{b_{n}})$
if they differ by a finite sequence of moves each of which either multiplies the coordinates by 1-1 or is of the form described in Remark 2.1(1). Equivalent sets of Seifert invariants yield homeomorphic manifolds.
4.1. $\mathcal{B}$ is spherical
When $\mathcal{B}$ is spherical, $n=3$ and $(b_{1},b_{2},b_{3})$ is one of the Platonic triples $(2,2,b_{3})$, $(2,3,3)$, $(2,3,4)$, or $(2,3,5)$. In each case, $b_{1}=2$ and $b_{2}$ is either $2$ or $3$. The fundamental group of $M$ is finite in all cases and therefore $b_{1}b_{2}b_{3}e(M)=\pm|H_{1}(M)|\neq 0$ (i.e., $H_{1}(M)$ is finite). Further, $\chi(\mathcal{B})>0$.
By assumption, $a_{i}=1$ when $b_{i}=2$. If $b_{2}=3$ and $a_{2}=2$, then the Seifert invariants of $M$ are
$(b,\frac{1}{2},\frac{2}{3},\frac{a_{3}}{b_{3}})\sim(b+1,\frac{1}{2},\frac{-1}{3},\frac{a_{3}}{b_{3}})\sim(-b-1,-\frac{1}{2},\frac{1}{3},-\frac{a_{3}}{b_{3}})\sim(-b-3,\frac{1}{2},\frac{1}{3},\frac{b_{3}-a_{3}}{b_{3}})$
Thus we can assume that $M$ has Seifert invariants $(b,\frac{1}{2},\frac{1}{b_{2}},\frac{a}{b_{3}})$. The following table lists $\chi(\mathcal{B}),e(M)$ and $|H_{1}(M)|$ for these manifolds. The last column determines the Seifert invariants of $M$ when $e(\nu_{M})=0$ by applying Theorem 1.1 and Corollary 1.3.
Seifert invariants $\chi(\mathcal{B})$ $e(M)$ $\pm|H_{1}(M)|$ Seifert invariants when $e(\nu_{M})=0$
$(b,\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2})$ $\tfrac{1}{2}$ $\tfrac{2b+3}{2}$ $4(2b+3)$ $(-1,\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2})$ and $(-2,\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2})$
$(b,\tfrac{1}{2},\tfrac{1}{2},\tfrac{a}{b_{3}}),\,b_{3}\geq 3$ $\tfrac{1}{b_{3}}$ $\tfrac{b_{3}(b+1)+a}{b_{3}}$ $4(b_{3}(b+1)+a)$ $(-1,\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{b_{3}})$
$(b,\tfrac{1}{2},\tfrac{1}{3},\tfrac{1}{3})$ $\tfrac{1}{6}$ $\tfrac{6b+7}{6}$ $3(6b+7)$ $(-1,\tfrac{1}{2},\tfrac{1}{3},\tfrac{1}{3})$
$(b,\tfrac{1}{2},\tfrac{1}{3},\tfrac{a}{4})$ $\tfrac{1}{12}$ $\tfrac{12b+10+3a}{12}$ $2(12b+10+3a)$ $(-1,\tfrac{1}{2},\tfrac{1}{3},\tfrac{1}{4})$
$(b,\tfrac{1}{2},\tfrac{1}{3},\tfrac{a}{5})$ $\tfrac{1}{30}$ $\tfrac{30b+6a+25}{30}$ $30b+6a+25$ $(-1,\tfrac{1}{2},\tfrac{1}{3},\tfrac{1}{5})$
It is interesting to note that the values of the Seifert invariants when $e(\nu_{M})=0$ correspond to those which realise the minimal values of $|H_{1}(M)|$ in the given families.
4.2. $\mathcal{B}$ is Euclidean
Here, either $n=3$ and $(b_{1},b_{2},b_{3})$ is one of the triples $(2,3,6),(2,4,4)$, $(3,3,3)$ or $n=4$ and $(b_{1},b_{2},b_{3},b_{4})=(2,2,2,2)$. We can assume that $a_{i}=1$ in all cases. This is obvious when $b_{i}=2$ and using Lemma 4.3 and the arguments of §4.1 deals with the case that some $b_{i}>2$.
As $\mathcal{B}$ is Euclidean, its Euler characteristic is zero. Hence if $\nu_{M}$ is trivial, Corollary 1.3 implies that $e(M)=0$ and we can use this to determine the Seifert invariants of $M$ when $e(\nu_{M})=0$.
Seifert invariants $e(M)$ $\pm|H_{1}(M)|$ Seifert invariants when $e(\nu_{M})=0$
$(b,\tfrac{1}{2},\tfrac{1}{3},\tfrac{1}{6})$ $b+1$ $36(b+1)$ $(-1,\tfrac{1}{2},\tfrac{1}{3},\tfrac{1}{6})$
$(b,\tfrac{1}{2},\tfrac{1}{4},\tfrac{1}{4})$ $b+1$ $32(b+1)$ $(-1,\tfrac{1}{2},\tfrac{1}{4},\tfrac{1}{4})$
$(b,\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3})$ $b+1$ $27(b+1)$ $(-1,\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3})$
$(b,\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2})$ $b+2$ $16(b+2)$ $(-2,\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2})$
As in the spherical case, those $M$ with $e(\nu_{M})=0$ are characterised by their first homology groups having minimal order (0 in this case). Their first Betti numbers are positive and therefore they admit horizontal fibrations (by tori) over the circle. This proves Theorem 1.7 when $\mathcal{B}$ is Euclidean.
4.3. $\mathcal{B}$ is hyperbolic
We assume in this section that the Seifert invariants of
$M=M(0;b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$
have been normalised so that $0<a_{i}<b_{i}$ for $1\leq i\leq n$.
The results of [12] and [15] show that $M$ admits a horizontal foliation if and only if one of the following conditions holds:
  1. (1)
    $2-n\leq b\leq-2$;
  2. (2)
    b=1b=-1 and there are coprime integers $0<c<d$ such that for some permutation $(\frac{c_{1}}{d},\frac{c_{2}}{d},\ldots,\frac{c_{n}}{d})$ of $(\frac{c}{d},\frac{d-c}{d},\frac{1}{d},\ldots,\frac{1}{d})$ we have $\frac{a_{i}}{b_{i}}<\frac{c_{i}}{d}$ for each $i$.
  3. (3)
    b=1nb=1-n and there are coprime integers $0<c<d$ such that for some permutation $(\frac{c_{1}}{d},\frac{c_{2}}{d},\ldots,\frac{c_{n}}{d})$ of $(\frac{c}{d},\frac{d-c}{d},\frac{1}{d},\ldots,\frac{1}{d})$ we have $\frac{b_{i}-a_{i}}{b_{i}}<\frac{c_{i}}{d}$ for each $i$.
The following specific examples illustrate that there are many Seifert manifolds $M$ that have co-orientable taut foliations with non-zero Euler class.
Examples 4.4.
Using Lemma 4.3, if some $b_{i},b_{j}$ have a common factor greater than $2$, one can construct many examples $M$ with $e(\nu_{M})\neq 0$. On the other hand using the conditions (1), (2), (3) immediately above, one can arrange for many of these to have horizontal foliations.
In (1), (2), (3), and (4) below we give some other classes of examples of Seifert fibred manifolds which admit co-orientable taut foliations but none with vanishing Euler class. The proofs are all applications of Theorem 1.1; we omit the details.
(1) By Lemma 4.3, if $b_{1}=b_{2}=\cdots=b_{n}$ and $e(\nu_{M})=0$, then $a_{1}=a_{2}=\cdots=a_{n}$. The first set of examples deals with a special case of this situation.
For $n\geq 4$, consider the manifold $M=M(0;b,\frac{1}{c},...,\frac{1}{c})$. Then
  • (a)
    $M$ has a horizontal foliation if and only if either $2-n\leq b\leq-2$, or b=1b=-1 and $c\geq 3$.
  • (b)
    If $c\geq 2n-2$, then $e(\nu_{M})=0$ if and only if b=2nb=2-n.
The bound on $c$ in part (b) is optimal: if c=2n3c=2n-3 and b=1b=-1, then $e(\nu_{M})=0$.
(2) For $n\geq 3$, let $M=M(0;2-n,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$, where $\sum_{i}\frac{a_{i}}{b_{i}}\leq\frac{n-2}{2}$. Then $M$ has a horizontal foliation, but $e(\nu_{M})=0$ if and only if $a_{i}=1$ for $1\leq i\leq n$.
(3) For $n\geq 4$, consider the manifold $M=M(0;b,\frac{1}{b_{1}},\ldots,\frac{1}{b_{n}})$, where the $b_{i}$ are pairwise coprime and $\geq n$. Then
  • (a)
    $M$ has a horizontal foliation if and only if $2-n\leq b\leq-1$.
  • (b)
    $e(\nu_{M})=0$ if and only if b=2nb=2-n.
(4) For $n\geq 4$, let $M=M(0;b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$, where $\sum_{i}\frac{a_{i}}{b_{i}}\leq\frac{n}{6}$. Then
  • (a)
    $M$ has a horizontal foliation for $2-n\leq b\leq-2$, and for b=1b=-1 if $n=4,5$, or $6$.
  • (b)
    $e(\nu_{M})=0$ for at most one $b$ with $2-n\leq b\leq\frac{2-n}{2}$.
4.4. Small values of $|H_{1}(M)|$
It is known that if $|H_{1}(M)|\leq 2$, then the Euler class $e(\nu_{M})$ vanishes (see [9, Proposition 2.1] for instance). In this section, we show that this fact can also be obtained from the computations in this article.
We have already discussed the case $M=M(0;b)$ in Remark 2.2, so we assume that $n>0$.
We recall the presentation of $H^{2}(\pi_{1}(\mathcal{B}))$ in §3.1, where $H^{2}(\pi_{1}(\mathcal{B}))$ is generated by $A_{0},A_{1},\ldots,A_{n}$ with relations $A_{0}=b_{i}A_{i}$ for $i=1,\ldots,n$. In §3, we showed that $e(\nu_{M})=\alpha^{*}(e(\overline{\varphi}))=0$, where $\overline{\varphi}\colon\pi_{1}(\mathcal{B})\to PSL(2,\mathbb{R})$ is discrete and faithful, if and only if $\exists\,m\in\mathbb{Z}$ such that $e(\overline{\varphi})=m\,e(E_{M})$. See (3.1.2) and (3.1.3). When $M=M\bigl(0;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, by (3.1.1) and (3.2.2),
$e(E_{M})=-\bigl(bA_{0}+a_{1}A_{1}+\cdots+a_{n}A_{n}\bigr),$
and
$e(\overline{\varphi})=-\left((2-n)A_{0}+A_{1}+A_{2}+\cdots+A_{n}\right)$
Suppose that $m$ is an integer such that $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i})$ for $i=1,\cdots,n$. Set
$S_{0}:=m(bA_{0}+a_{1}A_{1}+\cdots+a_{n}A_{n})=(mb+\sum_{i=1}^{n}\left\lfloor\frac{ma_{i}}{b_{i}}\right\rfloor)A_{0}+A_{1}+\cdots+A_{n}$
Any other integer $m^{\prime}$ for which the congruences $m^{\prime}a_{i}\equiv 1\,(\text{{\rm mod} }b_{i}),\text{ for }i=1,\cdots,n$ hold can be written $m^{\prime}=m+rl$, where $l$ is the least common multiple of $b_{1},b_{2},\ldots,b_{n}$ and $r\in\mathbb{Z}$. Set $m_{r}=m+rl$. The reader will verify that
(4.4.1) $S_{r}:=m_{r}(bA_{0}+a_{1}A_{1}+\cdots+a_{n}A_{n})=S_{0}+rle(M)A_{0}$
We noted after the statement of Corollary 1.3 that $e(M)=\varepsilon|H_{1}(M)|/b_{1}b_{2}\cdots b_{n}$ for some $\varepsilon\in\{\pm 1\}$. Thus from (4.4.1) we have
(4.4.2) $S_{r}=S_{0}+\varepsilon r\frac{|H_{1}(M)|}{(b_{1}b_{2}\cdots b_{n}/l)}A_{0}$
When $M$ is an integer homology $3$-sphere, $H^{2}(M)=0$, so $\nu_{M}$ is trivial. This is reflected in the fact that $l=b_{1}b_{2}\cdots b_{n}$ in this case, so we can choose $r$ so that the coefficient of $A_{0}$ in $S_{r}$ is 2n2-n.
Similarly, if $H_{1}(M)\cong\mathbb{Z}/2$, the fact that the tangent bundle to the Seifert fibres of $M$ is a trivial line bundle combines with the method of proof of [9, Proposition 2.1] to show that $e(\nu_{M})=0$. We leave it to the reader to verify this using (4.4.2) and Theorem 1.1.