### v2

# On $3$-manifolds admitting co-orientable taut foliations, but none with vanishing Euler class

Abstract.
In this article, we construct infinitely many (small Seifert fibred, hyperbolic and toroidal) rational homology $3$-spheres that admit co-orientable taut foliations, but none with vanishing Euler class. In the context of the $L$-space conjecture, these examples provide rational homology $3$-spheres that admit co-orientable taut foliations (and hence are not $L$-spaces) and have left-orderable fundamental groups, yet none of the left orders arise directly from the universal circle actions associated to co-orientable taut foliations.
The hyperbolic and non-Seifert toroidal examples are obtained from Dehn surgeries on knots in the $3$-sphere and use Heegaard Floer homology to obstruct the existence of a co-orientable foliation with vanishing Euler class. For the Seifert fibred case, we establish necessary and sufficient conditions for the Euler class of the normal bundle of the Seifert fibration to vanish. Moreover, when the base orbifold is hyperbolic, we also provide a second proof of this condition from the viewpoint of discrete faithful representations of Fuchsian groups.
Steven Boyer was partially supported by NSERC grant RGPIN 2024-06196
Ying Hu was partially supported by NSF grant DMS-2409398
Duncan McCoy was partially supported by NSERC grant RGPIN-2020-05491 and a Canada Research Chair
2010 Mathematics Subject Classification. Primary 57M25, 57M50, 57M99
Key words: taut foliation, Knot Floer complex, self-conjugate spinc structure, Seifert fibred manifold, normal bundle, Euler class, horizontal foliation.
1. Introduction
1.1. Rational homology spheres admitting no taut foliations with vanishing Euler class
It is predicted by the $L$-space conjecture [4, 18] that if a closed, orientable, prime $3$-manifold $M$ admits a co-orientable taut foliation, then its fundamental group is left-orderable. One of the main tools for proving this is to use Thurston’s universal circle action, which to each co-oriented taut foliation $\mathcal{F}$ on an oriented rational homology $3$-sphere $M$ associates a non-trivial representation
$\rho_{\mathcal{F}}:\pi_{1}(M)\to\mbox{Homeo}_{+}(S^{1}).$
See [10] for the details. If one can show that this representation lifts to the universal covering group $\mbox{Homeo}_{\mathbb{Z}}(\mathbb{R})$ of $\mbox{Homeo}_{+}(S^{1})$, then it follows from [8, Theorem 1.1] that the fundamental group of $M$ is left-orderable. However, such a lift exists if and only if the Euler class of $\rho_{\mathcal{F}}$ vanishes. Since this Euler class coincides with the Euler class of the tangent bundle of $\mathcal{F}$ [5], which we denote by $e(\mathcal{F})$, we are naturally led to the following question.
Question 1.1.
If a closed, orientable, prime $3$-manifold admits a co-orientable taut foliation, does it admit one whose Euler class is zero?  [^1] Given a co-orientable taut foliation, the Euler class of its tangent bundle is well-defined up to sign. So it makes sense to discuss if a co-orientable taut foliation has vanishing Euler class.
Although it has been expected that the answer to the question is no, we do not know of any examples verifying this. In this article, we show that such examples occur quite frequently.
Convention: When referring to $p/q$-surgery on a knot, we assume that $p$ and $q$ are coprime integers with $p\geq 0$.
Theorem 1.2.
Let $M$ be obtained by $p/q$-surgery on a knot $K$ in the $3$-sphere. If $q$ is even and |p/q|>2g(K)1|p/q|>2g(K)-1, then $M$ does not admit a co-orientable taut foliation with Euler class zero.
Thus the existence of a knot for which some large surgery with even denominator admits a co-orientable taut foliation shows that the answer to Question 1.1 is no. For instance, Roberts has shown that if $K$ is a non-trivial fibred knot in the $3$-sphere, then $K(p/q)$, the manifold obtained by $p/q$-surgery on $K$, admits a co-orientable taut foliation for all $p/q$, $p/q<1$, or p/q>1p/q>-1 depending on whether the fractional Dehn twist coefficient of $K$ is zero, positive, or negative respectively [26]. The following corollary states the case when the fractional Dehn twist coefficient is zero. Examples of such knots include all hyperbolic fibred alternating knots [6, 2].
Corollary 1.3.
Suppose that $K$ is a non-trivial fibred knot in the $3$-sphere whose fractional Dehn twist coefficient is zero. Then for all slopes $p/q$ with $q$ even and |p/q|>2g(K)1|p/q|>2g(K)-1, $K(p/q)$ admits a co-orientable taut foliation, but none with zero Euler class.
A similar conclusion holds for many non-fibred knots. Indeed, the $L$-space conjecture predicts that every irreducible rational surgery on a non-fibred knot admits a co-orientable taut foliation and this is known to be true in many instances. For instance, Roberts and Delman proved this for non-torus alternating knots and for Montesinos knots which are not $L$-space knots.
Examples of manifolds having left-orderable fundamental groups and admitting co-orientable taut foliations, but none with vanishing Euler class, can also be constructed, thus showing that no left-orders on the group arise from lifts of universal circle representations. For instance, the figure eight knot is a fibred hyperbolic knot of genus one with zero fractional Dehn twist coefficient and each rational surgery on it has a left-orderable fundamental group [27].
Corollary 1.4.
Suppose that $K$ is the figure eight knot. Then for all slopes $p/q$ with $q$ even and $|p/q|>1$, $K(p/q)$ has a left-orderable fundamental group and admits a co-orientable taut foliation, but no co-orientable taut foliation with zero Euler class.
By way of contrast, we note that Theorem 1.2 cannot be extended to all large surgeries on a knot without some hypotheses on the slope $p/q$. For instance, it is known that $p$-surgery on the figure eight knot admits co-orientable taut foliations of zero Euler class for each integer $p$ [11, 14]. More generally, we have
Proposition 2.3.
Let $M$ be a rational homology sphere obtained by integer surgery on a genus one knot in the $3$-sphere. Then every co-orientable taut foliation on $M$ has Euler class zero.
Remark 1.5.
The line of reasoning used in the proof of Proposition 2.3 can easily be extended to show that if $K$ is an almost $L$-space knot, meaning a knot with positive almost $L$-space surgeries [9], and $M$ is obtained by $p$-surgery on $K$ for some $p\geq 2g(K)-1$, then any co-orientable taut foliation on $M$ has Euler class zero. This shows that for knots of arbitrarily large genus, constraints must be placed on the denominators of surgery slopes for Theorem 1.2 to hold.
Corollary 1.6.
Suppose that $K$ is a knot of genus one in the $3$-sphere. If $p$ is an integer such that $K(p)$ admits a co-orientable taut foliation, then $\pi_{1}(K(p))$ is left-orderable. Similarly, if $K$ is an almost $L$-space knot and $p\geq 2g(K)-1$ is an integer such that $K(p)$ admits a co-orientable taut foliation, then $\pi_{1}(K(p))$ is left-orderable.
All the examples in Corollary 1.4 are hyperbolic. The simplest examples of Seifert fibred manifolds in which every co-orientable taut foliation has non-zero Euler class arise from Dehn surgery on the trefoil knot.
Theorem 1.7.
Suppose that $K$ is the right-handed trefoil knot. Then $K(p/q)$ admits a co-orientable taut foliation and has a left-orderable fundamental group if and only if $p/q<1$. It admits a co-orientable taut foliation with zero Euler class if and only if $p/q<1$ and $|q|\equiv 1\pmod{p}$.
In general, we denote a closed oriented Seifert fibred manifold $M$ with orientable base orbifold by $M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, where $b\in\mathbb{Z}$ and $b_{i}\geq 2$ for $i=1,\ldots,n$. See §3.1 for a detailed explanation of the notation. We denote the base orbifold by $\mathcal{B}$; this is a closed orientable surface $|\mathcal{B}|$ of genus $g\geq 0$ with $n\geq 0$ cone points of orders $b_{1},...,b_{n}$. Recall that the Euler characteristic $\chi(\mathcal{B})$ of $\mathcal{B}$ is $\chi(|\mathcal{B}|)-n+\sum_{i=1}^{n}\frac{1}{b_{i}}$, and the Euler number of $M$ is $e(M)=b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}$.
Theorem 1.8.
Suppose that $M=M\bigl(0;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ is a rational homology $3$-sphere which admits a co-orientable taut foliation $\mathcal{F}$. Then $\mathcal{F}$ has zero Euler class if and only if there exists $m\in\mathbb{Z}$ such that
  1. (1)
    $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i}),\text{ for }i=1,\cdots,n,\;\text{and}$
  2. (2)
    $me(M)=\chi(\mathcal{B})$.
We can also produce toroidal examples which are not Seifert fibred. Define the JSJ graph of a knot $K$ to be the rooted tree dual to the JSJ tori of the exterior $X(K)$ of $K$, where the root vertex corresponds to the piece of the JSJ decomposition containing $\partial X(K)$. Any finite rooted tree can arise as the JSJ graph of a knot. We say that a JSJ graph is a rooted interval when it is homeomorphic to an interval with root at an endpoint. Since non-integral surgeries on knots in $S^{3}$ are irreducible [12], the following corollary follows immediately from Theorem 1.2 and Theorems 3.13 and 4.3 of [3].
Corollary 1.9.
Suppose that $K$ is a fibred satellite knot whose $JSJ$ graph is not a rooted interval. Then for all slopes $p/q$ with $q$ even and |p/q|>2g(K)1|p/q|>2g(K)-1, $K(p/q)$ is toroidal, has a left-orderable fundamental group, and admits a co-orientable taut foliation, but none with zero Euler class.
Remark 1.10.
Fibred composite knots satisfy the corollary’s hypothesis. Taking $K$ to be a connected sum of torus knots yields examples which are graph manifolds.
1.2. Euler class of the normal bundle of a Seifert fibration
While Theorem 1.2 is proved using the knot Floer complex, the key fact in analyzing the Euler classes of co-oriented taut foliations on a Seifert fibred rational homology $3$-sphere $M$ is that, up to isotopy, all such foliations are horizontal; that is, transverse to the Seifert fibres of $M$ (Lemma 5.1). Hence, the tangent bundle of $\mathcal{F}$ is isomorphic to the normal bundle $\nu_{M}$ to the Seifert fibres of $M$, while the latter makes sense even without the existence of a horizontal foliation.
We will refer to $\nu_{M}$ as the normal bundle of the Seifert fibration. In §3 and §4, we determine the following necessary and sufficient conditions for its Euler class to vanish.
Theorem 1.11.
Given a closed oriented Seifert fibred manifold $M=M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, the Euler class of the normal bundle $\nu_{M}$ of the Seifert fibration vanishes if and only if there exists $m\in\mathbb{Z}$ such that
  1. (1)
    $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i}),\text{ for }i=1,\cdots,n,\;\text{and}$
  2. (2)
    $me(M)=\chi(\mathcal{B})$.
Theorem 1.8 follows from Theorem 1.11 and Lemma 5.1.
Next, we deduce the following easily verifiable necessary condition for $\nu_{M}$ to be a trivial bundle:
Corollary 1.12.
Given a closed oriented Seifert fibred manifold $M$, if the Euler class of the normal bundle $\nu_{M}$ of the Seifert fibration vanishes, then either $M$ is an orientable $S^{1}$-bundle over the torus or one of the following holds:
  1. (1)
    $e(M)=\chi(\mathcal{B})=0$,
  2. (2)
    $e(M)\neq 0,\,\chi(\mathcal{B})\neq 0,$ and
    $\frac{\chi(\mathcal{B})}{e(M)}\in\mathbb{Z}.$
Note that if $M$ is an orientable $S^{1}$-bundle over the torus, then its normal bundle is always trivial (see Remark 3.2).
By abelianizing the presentation (4.0.1) of $\pi_{1}(M)$ obtained by viewing $M$ as doing a Dehn filling of a trivial $S^{1}$-bundle over an $n$-punctured orientable surface of genus $g$ as in §3.1, one can show that if $T_{1}(M)$ denotes the torsion subgroup of $H_{1}(M)$, then
$|T_{1}(M)|=b_{1}\cdots b_{n}|e(M)|,$
where we take the convention that $b_{1}b_{2}\cdots b_{n}=1$ when $n=0$. Thus, it follows from Corollary 1.12 that:
Corollary 1.13.
Consider the Seifert fibred manifold $M=M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ and suppose that the Euler class of the normal bundle $\nu_{M}$ of the Seifert fibration vanishes. Then the integer $b_{1}\ldots b_{n}\chi(\mathcal{B})$ is a multiple of $|T_{1}(M)|$. Hence, if $M$ is a rational homology $3$-sphere, then $b_{1}\ldots b_{n}\chi(\mathcal{B})$ is a multiple of $|H_{1}(M)|$.
We will present more examples of rational homology Seifert fibred manifolds that admit co-orientable taut foliations, but none with vanishing Euler class, in §5.3 (Examples 5.3). In fact, our computations in §5 suggest that most closed Seifert fibred manifolds have nontrivial normal bundle, including the case when the base orbifold is Euclidean (i.e. $\chi(\mathcal{B})=0$). This may seem counterintuitive, since we have pointed out that the normal bundle is always trivial when the base $\mathcal{B}$ is the torus.
We close with an interesting by-product of our analysis.
Theorem 1.14.
Let $M$ be a closed oriented Seifert fibred manifold with orientable base orbifold $\mathcal{B}$. If $\mathcal{B}$ is hyperbolic or Euclidean, and $e(\nu_{M})=0$, then $M$ admits a horizontal foliation.
In contrast, closed Seifert fibred manifolds with spherical base orbifolds have finite fundamental groups and hence cannot admit taut foliations. Yet, in §5.1, we will show examples of such manifolds for which $e(\nu_{M})$ vanishes.
Plan of the paper
We prove Theorem 1.2 and Proposition 2.3 in §2. In §3 we explain our notation for Seifert manifolds and prove Theorem 1.11, which gives necessary and sufficient conditions for the Euler class of the normal bundle of a Seifert fibration to vanish. We also prove Corollary 1.12 in this section. In §4 we present a different perspective on Theorem 1.11 and Corollary 1.12 in the case that the base orbifold is hyperbolic through the use of discrete faithful $PSL(2,\mathbb{R})$-representations of $\pi_{1}(\mathcal{B})$. As a consequence of this approach, we prove the hyperbolic case of Theorem 1.14. Finally, in §5 we prove Theorems 1.7 and 1.8, and provide examples of Seifert fibred manifolds whose normal bundles are (non-)trivial. We classify which occurs when $\mathcal{B}$ is spherical in §5.1 and is Euclidean in §5.2. The Euclidean case of Theorem 1.14 is dealt with at the end of §5.2. The case that $\mathcal{B}$ is hyperbolic is analysed in §5.3.
Acknowledgements
We thank Nathan Dunfield for bringing up the question of finding examples of rational homology spheres that admit co-oriented taut foliations but none with vanishing Euler class, during his lecture at the workshop “Low Dimensional Topology and Floer Theory” in Montreal, August 18-29, 2025, which inspired us to investigate the Euler class of horizontal foliations in Seifert fibred manifolds. The second and third named authors would like to thank the Simons Foundation, the CRM, and CIRGET for their support during the preparation of this article.
2. Examples from surgeries on knots in the $3$-sphere
Theorem 1.2 is an immediate consequence of the following two lemmas. The first follows from Ozsváth and Szabó’s proof that a manifold with a co-orientable taut foliation is not an $L$-space [22].
Lemma 2.1.
Let $M$ be a rational homology $3$-sphere admitting a co-orientable taut foliation $\mathcal{F}$. Then $M$ has a $\mathrm{spin}^{c}$-structure $\mathfrak{s}$ such that $HF_{\mathrm{red}}(M,\mathfrak{s})\neq 0$ and $c_{1}(\mathfrak{s})=e(\mathcal{F})$.
Proof.
Let $M$ be a rational homology $3$-sphere that admits a co-orientable taut foliation $\mathcal{F}$. The tangent plane field of $\mathcal{F}$ can be perturbed to obtain a tight contact structure $\xi$ whose contact invariant $c(\xi)\neq 0\in HF_{\mathrm{red}}(M,\mathfrak{s})$, with $c_{1}(\mathfrak{s})=e(\mathcal{F})$ [22]. ∎
The second is an application of the mapping cone formula for calculating the Heegaard Floer homology of surgeries on a knot in the $3$-sphere [23, 24]. In conjunction with Lemma 2.1, this implies Theorem 1.2 since for a 3-manifold $M$ for which $H_{1}(M)$ has odd order, the unique spinc-structure satisfying $c_{1}(\mathfrak{s})=0$ is also the unique self-conjugate spinc-structure on $M$.
Lemma 2.2.
Let $M$ be obtained by $p/q$-surgery on a knot $K$ in the $3$-sphere, where $q$ is even and |p/q|>2g(K)1|p/q|>2g(K)-1. Then the unique self-conjugate $\mathrm{spin}^{c}$-structure $\mathfrak{s}$ on $M$ satisfies $HF_{\mathrm{red}}(M,\mathfrak{s})=0$.
Proof.
Without loss of generality we may assume that $p/q>0$. Let $CFK^{\infty}(K)=\bigoplus_{i,j\in\mathbb{Z}}C\{(i,j)\}$ be the knot Floer complex for $K$. Since |p/q|>2g(K)1|p/q|>2g(K)-1, the mapping cone formula for computing $HF^{+}(M)$ in terms of $CFK^{\infty}(K)$ implies that there is a labelling
$\phi:\mathbb{Z}/p\to\mathrm{Spin}^{c}(M),$
such that [21, Proposition 3.5]
$HF_{\mathrm{red}}(M,\phi(i))\cong\bigoplus_{k\in\mathbb{Z}}\bm{A}_{{\rm red},\,\lfloor\frac{pk+i}{q}\rfloor}(K),$
where $\bm{A}_{{\rm red},\,k}(K)$ denotes the reduced part of
$\bm{A}^{+}_{k}=H_{*}(C\{i\geq 0\text{ or }j\geq k\}).$
However, $\bm{A}_{{\rm red},\,k}(K)=0$ for |k|>g(K)1|k|>g(K)-1 [21, Lemma 3.2]. Thus, for $0\leq i\leq p-1$, this becomes
$HF_{\mathrm{red}}(M,\phi(i))\cong\begin{cases}\bm{A}_{{\rm red},\,\lfloor\frac{i}{q}\rfloor}(K),&0\leq i<qg(K)-q,\\ \bm{A}_{{\rm red},\,\lfloor\frac{i-p}{q}\rfloor}(K),&p-qg(K)<i<p,\\ 0,&qg(K)-q\leq i\leq p-qg(K).\end{cases}$
In particular, for each $|j|\leq g(K)-1$, there are precisely $q$ $\mathrm{spin}^{c}$-structures for which $HF_{\mathrm{red}}(M,\mathfrak{s})$ is isomorphic to $\bm{A}_{{\rm red},\,j}(K)$, and for all other $\mathrm{spin}^{c}$-structures, $HF_{\mathrm{red}}(M,\mathfrak{s})=0$. Thus the set
$S=\{\mathfrak{s}\in\mathrm{Spin}^{c}(M)\mid HF_{\mathrm{red}}(M,\mathfrak{s})\neq 0\}$
has cardinality divisible by $q$. In particular, $|S|$ is even. Since conjugation of $\mathrm{spin}^{c}$-structures induces an isomorphism of Heegaard Floer homology, conjugation acts as an involution on $S$. This implies that there are an even number of self-conjugate $\mathrm{spin}^{c}$-structures in $S$. Since $p$ is odd, there is precisely one self-conjugate spinc-structure on $M$, so it follows that $S$ does not contain the self-conjugate spinc-structure on MM. [^2] An alternative line of argument is to use the fact that conjugation of spinc-structures corresponds to the involution $i\mapsto p+q-1-i$ (mod $p$) on $\mathbb{Z}/p$. For $q$ even, this implies that the self-conjugate spinc-structure is labelled by $s=\frac{p+q-1}{2}$. However since |p/q|>2g(K)1|p/q|>2g(K)-1, the above calculation shows that $HF_{\mathrm{red}}(M,\phi(s))=0$.
In the opposite direction, we note that Theorem 1.2 cannot be extended to all large surgeries on a knot $K$ without some hypotheses on the slope $p/q$.
Proposition 2.3.
Let $M$ be a rational homology sphere obtained by integer surgery on a genus one knot in the 3-sphere. Then every co-orientable taut foliation on $M$ has Euler class zero.
Proof.
Let $K$ be a genus one knot and let $M$ be obtained by $p$-surgery on $K$. Without loss of generality we may assume that $p>0$. The proposition is deduced from Lemma 2.1 by showing that there is at most one spinc-structure $\mathfrak{s}_{0}$ for which $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})\neq 0$ and then verifying that $c_{1}(\mathfrak{s}_{0})=0$.
First we compute $HF_{\mathrm{red}}(M)$. For integer surgeries of slope $p\geq 2g(K)-1$, the mapping cone formula gives a labelling
(2.0.1) $\phi:\mathbb{Z}/p\to\mathrm{Spin}^{c}(M),$
such that [21, Proposition 3.5]
$HF_{\mathrm{red}}(M,\phi(i))\cong\bigoplus_{k\in\mathbb{Z}}\bm{A}_{{\rm red},\,i+kp}(K).$
However, as $g(K)=1$, we have $\bm{A}_{{\rm red},\,k}(K)=0$ for $|k|>0$ [21, Lemma 3.2]. Thus, for $0\leq i\leq p-1$, this becomes
$HF_{\mathrm{red}}(M,\phi(i))\cong\begin{cases}\bm{A}_{{\rm red},\,0}(K),&i=0,\\ 0,&0<i\leq p-1,\end{cases}$
giving that that there is at most one spinc-structure $\mathfrak{s}_{0}$ for which $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})\neq 0$ and that this is the spinc-structure corresponding to 0 under (2.0.1). We verify that $c_{1}(\mathfrak{s}_{0})=0$ by unpacking the labelling in (2.0.1). As in [23, Section 2.4] the map $\phi$ is obtained as follows. Take $X$ to be the trace of $p$-surgery on $K$. Let $\Sigma]\in H_{2}(X)$ denote that the homology class obtained by capping off the core of the 2-handle with a Seifert surface for $K$. If $\mathfrak{s}\in\mathrm{Spin}^{c}(M)$ is the restriction of $\mathfrak{t}\in\mathrm{Spin}^{c}(X)$, then $\phi(i)=\mathfrak{s}$, where $i$ satisfies
$\langle c_{1}(\mathfrak{t}),[\Sigma]\rangle+p\equiv 2i\bmod 2p.$
Thus the spinc-structure with $\mathfrak{s}_{0}=\phi(0)$ is the restriction of any spinc-structure $\mathfrak{t}_{0}$ on $X$ such that $\langle c_{1}(\mathfrak{t}_{0}),[\Sigma]\rangle\equiv p\bmod 2p$. Since the restriction map $H^{2}(X)\rightarrow H^{2}(M)$ is reduction mod $p$, this implies that $c_{1}(\mathfrak{s}_{0})=0$, as required. ∎
Remark 2.4.
As mentioned in the introduction, the line of reasoning in the proof of Proposition 2.3 can easily be extended to show that if $K$ is an almost $L$-space knot and $M$ is obtained by $p$-surgery on $K$ for some $p\geq 2g(K)-1$, then any co-orientable taut foliation on $M$ has Euler class zero.
3. The Euler class of the normal bundle of a Seifert fibration
Let $M$ be a closed, oriented Seifert fibred manifold whose base orbifold has an orientable underlying space. In this section, we compute the Euler class of the normal bundle of the Seifert fibration on $M$ and determine conditions for it to vanish (Theorem 1.11 and Corollary 1.12).
3.1. Seifert notation
We first recall the Seifert notation that we will use to describe $M$.
Consider $M$ constructed by attaching $n+1$ solid tori $N_{i}$, $i=0,1,\dots,n$, to a trivial $S^{1}$-bundle over an orientable surface $F$ with $|\partial F|=n+1$, where we assume that $N_{0}$ is a tubular neighbourhood of a regular fibre and the $N_{i}$ are tubular neighbourhoods of the singular fibres for $i=1,2,\ldots,n$. Denote this trivial $S^{1}$-bundle by $P\cong F\times S^{1}$. Fix orientations on $F$ and on the fibres so that they intersect in $M$.
Let
$\partial F=\mu_{0}\sqcup\mu_{1}\sqcup\cdots\sqcup\mu_{n}$
with the induced orientation, and
$\partial P=T_{0}\sqcup T_{1}\sqcup\cdots\sqcup T_{n}$
with $\mu_{i}\subset T_{i}$ for each $i=0,1,\cdots,n$. So on each $T_{i}$, we use the basis given by $\{h_{i},\mu_{i}\}$, where $h_{i}$ is given by a Seifert fibre. By our orientation convention, we have $\mu_{i}\cdot h_{i}=1$.
Let $D_{i}$ be a meridional disk of $N_{i}$, oriented so that it intersects the Seifert fibres positively. Under the attaching map, we have $\partial D_{i}=-\bigl(a_{i}h_{i}+b_{i}\mu_{i}\bigr)$ as oriented curves, for some $a_{i}\in\mathbb{Z}$, $b_{i}\in\mathbb{Z}_{>0}$ and $a_{i}$, $b_{i}$ coprime. In fact, since $N_{0}$ is a tubular neighbourhood of a regular fibre, and the $N_{i}$ are tubular neighbourhoods of singular fibres, we have $b_{0}=1$ and $b_{i}\geq 2$.
Let $g\geq 0$ denote the genus of $F$, $b=\frac{a_{0}}{b_{0}}\in\mathbb{Z}$, and denote
$M=M\Bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\Bigr)$
Let $\mathcal{B}$ denote the base orbifold of $M$, which is a genus $g$ closed orientable surface with $n$ cone points of orders $b_{1},\ldots,b_{n}$ respectively. We use $|\mathcal{B}|$ to denote the underlying space of $\mathcal{B}$.
Remarks 3.1.
  1. (1)
    This description of the Seifert fibred manifold is not unique. More precisely, we have that $M\Bigl(g;\,\tfrac{a_{0}}{b_{0}},\ldots,\tfrac{a_{n}}{b_{n}}\Bigr)$ is homeomorphic to $M\Bigl(g;\,\tfrac{a^{\prime}_{0}}{b_{0}},\ldots,\tfrac{a^{\prime}_{n}}{b_{n}}\Bigr)$ by an orientation-preserving homeomorphism preserving fibres if and only if up to reordering,
    $\frac{a_{i}}{b_{i}}-\frac{a^{\prime}_{i}}{b_{i}}\in\mathbb{Z}\quad\text{and}\quad\sum_{i=0}^{n}\frac{a_{i}}{b_{i}}=\sum_{i=0}^{n}\frac{a^{\prime}_{i}}{b_{i}}.$
    By choosing exactly one of the $N_{i}$, namely $N_{0}$, to be a neighbourhood of a regular fibre in the construction above, we are adopting the usual convention of having only one of the $\tfrac{a_{i}}{b_{i}}$ be an integer.
  2. (2)
    Given $M=M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, the sum
    $e(M):=b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}$
    is an invariant of the given Seifert fibring of $M$ called the Euler number.
  3. (3)
    The effect on $b,a_{1},\ldots,a_{n}$ and $e(M)$ of reversing the orientation of $M$ is to replace them by their negatives.
Remark 3.2.
When $n=0$, the manifold $M=M(g;b)$ is an $S^{1}$-bundle over a closed, orientable surface $B$ of genus $g$ with Euler number $b$. Moreover, the necessary and sufficient condition in Theorem 1.11 simplifies to the existence of $m\in\mathbb{Z}$ such that
(3.1.1) $mb=\chi(B)$
By assumption, the Euler class of the circle bundle $f:M\to B$ is $bu$, where $u$ is a generator of $H^{2}(B)\cong\mathbb{Z}$. It is easy to see that $\nu_{M}$ is the pull-back by $f$ of the tangent bundle $\tau_{B}$ of $B$. Hence,
(3.1.2) $e(\nu_{M})=f^{*}(e(\tau_{B})),$
Let $\langle h\rangle$ denote the central subgroup of $\pi_{1}(M)$ generated by a circle fibre of $f$ and consider the $5$-term exact sequence in cohomology associated to the central extension $1\to\langle h\rangle\to\pi_{1}(M)\to\pi_{1}(B)\to 1$:
$0\to H^{1}(B)\xrightarrow{}H^{1}(M)\to\mathbb{Z}\xrightarrow{\;\delta\;}H^{2}(B)\xrightarrow{\;f^{*}\;}H^{2}(M)$
It follows from [13, Theorem 4] that the image of 1-1 under the transgression homomorphism $\delta$ is the Euler class $bu$ of the circle bundle. Hence $\mbox{kernel}(f^{*})$ is generated by $bu$. Then by (3.1.2), $e(\nu_{M})=0$ if and only $e(\tau_{B})=\pm\chi(B)u$ is a multiple of $bu$, which is condition (3.1.1).
We consider the following two special cases:
  1. (1)
    If $M=M(g;b)$ is a rational homology sphere, then the base $B$ must be the $2$-sphere, and $M=M(0;b)$ is a lens space $L(|b|,1)$. In this case, the Euler class $e(\nu_{M})$ vanishes if and only if $b\in\{\pm 1,\pm 2\}$.
  2. (2)
    When $B$ is the torus, (3.1.2) implies that the Euler class $e(\nu_{M})$ is always trivial, since the tangent bundle $T\mathcal{B}$ is trivial,
3.2. Proof of Corollary 1.12
We prove Corollary 1.12 using Theorem 1.11 here and defer the proof of Theorem 1.11 to the next subsection.
Let $\nu_{M}$ denote the normal plane bundle over $M$ to the Seifert fibration, oriented with the induced orientation from those on $M$ and on the fibres.
Proof of Corollary 1.12.
Assume that the Euler class of the normal bundle vanishes.
We first consider the case $n=0$, in which $\mathcal{B}$ is a genus $g$ surface and $M$ is an $S^{1}$-bundle over $\mathcal{B}$ of Euler number $b$. The necessary and sufficient condition in Theorem 1.11 is equivalent to that the existence of $m\in\mathbb{Z}$ such that $mb=\chi(\mathcal{B})$.
If $m=0$, then $\chi(\mathcal{B})=mb=0$ and therefore $M$ is an oriented $S^{1}$-bundle over the torus. On the other hand, if $M$ is not an oriented $S^{1}$-bundle over the torus, then $\chi(\mathcal{B})\neq 0$. Then for the equality $mb=\chi(\mathcal{B})$ to hold, we must have $e(M)=b\neq 0$. Hence
$m=\frac{\chi(\mathcal{B})}{e(M)}\in\mathbb{Z}.$
Next we consider the case $n>0$. Again by Theorem 1.11, there exists $m\in\mathbb{Z}$ such that $ma_{i}\equiv 1\pmod{b_{i}}$ and $me(M)=\chi(\mathcal{B})$. Since $m$ cannot be zero in this case, either (1) or (2) holds. ∎
3.3. Proof of Theorem 1.11
Here we provide a proof of Theorem 1.11. When the base orbifold is hyperbolic, i.e., $\chi(\mathcal{B})<0$, we will give another perspective on Theorem 1.11 and Corollary 1.12 in §4. There we will show that the Euler class of the normal bundle is precisely the pullback of the Euler class of a discrete faithful representation of $\pi_{1}(\mathcal{B})$ in $PSL(2,\mathbb{R})$ under the homomorphism $\pi_{1}(M)\to\pi_{1}(\mathcal{B})$ induced by the natural projection $M\rightarrow\mathcal{B}$  [^3] Since $M$ is irreducible when $\mathcal{B}$ is hyperbolic, we have $H^{2}(M)\equiv H^{2}(\pi_{1}(M))$..
Proof of Theorem 1.11.
To simplify the notation in the proof, we will use $\tfrac{a_{0}}{b_{0}}$ in place of $b$, as explained in §3.1. Thus
$M=M\bigl(0;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)=M\bigl(0;\,\tfrac{a_{0}}{b_{0}},\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$
We will compute the Euler class of the normal bundle, denoted by $e(\nu_{M})\in H^{2}(M)$, using the following exact sequence of the pair $(M,\sqcup_{i}T_{i})$:
$\cdots\to\oplus_{i}H^{1}(T_{i})\xrightarrow{\delta}H^{2}(M,\sqcup_{i}T_{i})\xrightarrow{i^{*}}H^{2}(M)\to 0$
Note that
$H^{2}(M,\sqcup_{i}T_{i})\cong H^{2}(P,\sqcup_{i}T_{i})\oplus H^{2}(N_{1},T_{1})\oplus\cdots\oplus H^{2}(N_{n},T_{n}),$
and by the Künneth formula,
$H^{2}(P,\sqcup_{i}T_{i})\cong H^{1}(F,\partial F)\otimes H^{1}(S^{1})\oplus H^{2}(F,\partial F)$
Recall that $\partial F=\sqcup_{i}\mu_{i}$ with $\mu_{i}\subset T_{i}$ for each $i=0,1,\ldots,n$. Suppose that oriented simple closed curves $\alpha_{r},\beta_{r}$, $r=1,\ldots,g$, together with oriented arcs $c_{j}$ on $F$ that connect the boundary component $\mu_{j}$ to the boundary component $\mu_{j+1}$, $j=0,1,\ldots,n-1$ when $n>0$, represent a set of generators of $H_{1}(F,\partial F)$. Then $H_{2}(P,\sqcup_{i}T_{i})$ is freely generated by the classes of vertical tori $\alpha_{r}\times S^{1}$, $\beta_{r}\times S^{1}$, the horizontal surface $F$ and vertical annuli $c_{j}\times S^{1}$ when $n>0$. We denote their duals in $H^{2}(P,\partial P)$ by
$\alpha_{r}^{*}\otimes S^{*},\quad\beta_{r}^{*}\otimes S^{*},\quad c_{j}^{*}\otimes S^{*},\quad F^{\ast}$
respectively, where $r=1,\ldots,g$ and $j=0,1,\ldots,n-1$.
Finally, let $D_{i}^{*}\in H^{2}(N_{i},T_{i})$ denote the dual of $D_{i},\partial D_{i}]\in H_{2}(N_{i},T_{i})$ for each $i=0,1,\ldots,n$.
In summary, a cohomology class in
$H^{2}(M,\sqcup T_{i})\cong H^{1}(F,\partial F)\otimes H^{1}(S^{1})\oplus H^{2}(F,\partial F)\oplus H^{2}(N_{1},T_{1})\oplus\cdots\oplus H^{2}(N_{n},T_{n})$
can be uniquely expressed as a linear combination of
$\alpha_{r}^{*}\otimes S^{*},\quad\beta_{r}^{*}\otimes S^{*},\quad F^{*},\quad D_{i}^{*}\quad\text{and}\quad c_{j}^{*}\otimes S^{*}\text{ when }n>0$
where $r=1,\ldots,g$, $j=0,1,\ldots,n-1$ and $i=0,1,\ldots,n$.
The proof proceeds in two steps (1) and (2).
  1. (1)
    Find the preimage of the Euler class $e(\nu_{M})$ of the normal bundle $\nu_{M}$ in $H^{2}(M,\sqcup T_{i})$.
    Let $\sigma$ be a nowhere vanishing vector field along $\partial P=\sqcup T_{i}$ that points out of $P$. Since $T_{i}$ is tangent to the Seifert fibres, $\sigma$ is a section of $\nu_{M}$ along $\sqcup T_{i}$. We give a cellular structure on $M$ for which we assume that the vertical surfaces $\alpha_{r}\times S^{1}$, $\beta_{r}\times S^{1}$, $c_{j}\times S^{1}$, and the horizontal surfaces $F$ and $D_{i}$, as well as the $T_{i}$, all belong to the 2-skeleton. We also assume that the curves $\alpha_{r}$, $\beta_{r}$ and the arcs $c_{j}$ on $F$ are in the 1-skeleton.
    From $\sigma$ one can define an obstruction $2$-cochain, denoted by $c_{\sigma}$, that represents the relative Euler class of the normal bundle $\nu_{M}$ in $H^{2}(M,\sqcup T_{i})$. By construction, the image of $c_{\sigma}]$ under $H^{2}(M,\sqcup T_{i})\to H^{2}(M)$ is the Euler class $e(\nu_{M})$. (See [14, Section 2] for instance.)
    To write $c_{\sigma}]$ as a linear combination of $\alpha_{r}^{*}\otimes S^{*}$, $\beta_{r}^{*}\otimes S^{*}$, $c_{j}^{*}\otimes S^{*}$, $F^{*}$, and $D_{i}^{*}$, it suffices to compute its evaluation on the dual homology classes.
    1. (a)
      $c_{\sigma}]$ evaluates to zero on all vertical surfaces. This is because, starting from $\sigma|_{\partial F}$, where $\partial F\subset\partial P=\sqcup T_{i}$, there is no obstruction to extending $\sigma$ to the 1-skeleton of $F$. Since $P\cong F\times S^{1}$ is a trivial $S^{1}$-bundle, one can extend the section further to the vertical surfaces $\alpha_{r}\times S^{1}$, $\beta_{r}\times S^{1}$ and $c_{j}\times S^{1}$ by staying constant along the fibre direction.
    2. (b)
      By the Poincaré-Hopf Theorem, $c_{\sigma}]$ evaluated on a horizontal surface equals its Euler characteristic, i.e.,
      $c_{\sigma}]([D_{i},\partial D_{i}])=1,\quad i=0,1,\ldots,n,$
      and
      $c_{\sigma}]([F,\partial F])=\chi(F)=\chi(|\mathcal{B}|)-(n+1).$
  2. (2)
    Compute the image of $\delta:\oplus_{i}H^{1}(T_{i})\rightarrow H^{2}(M,\sqcup_{i}T_{i})$.
    Let $h_{i}^{*},\mu_{i}^{*}\in H^{1}(T_{i})$ denote the (Hom) dual classes of $h_{i}],[\mu_{i}]\in H_{1}(T_{i})$ for $i=0,1,\ldots,n$. We compute $\delta h_{i}^{*}$ and $\delta\mu_{i}^{*}$, which generate the image of $\delta$.
    1. (a)
      It is easy to see that, since $\delta h_{i}^{*}=h_{i}^{*}\circ\partial$ and $\delta\mu_{i}^{*}=\mu_{i}^{*}\circ\partial$, they vanish on all vertical tori, $\alpha_{r}\times S^{1}$ and $\beta_{r}\times S^{1}$.
    2. (b)
      As $\partial([c_{j}\times S^{1}])=[h_{j+1}]-[h_{j}]$ in $H_{1}(\sqcup_{i}T_{i})$ for $j=0,1,\ldots,n-1$, we have
      $\delta\mu_{i}^{*}([c_{j}\times S^{1}])=0\quad\text{for all }i=0,1,\ldots,n,\text{ and }\ j=0,1,\ldots,n-1,$
      and
      $\delta h_{i}^{*}([c_{j}\times S^{1}])=\begin{cases}1&\text{if }j=i-1\\[6.0pt] -1&\text{if }j=i\\[6.0pt] 0&\text{otherwise}\end{cases}$
    3. (c)
      Since $\partial F=\sqcup_{i}\mu_{i}$, we have $\delta h_{i}^{*}([F,\partial F])=0$ and $\delta\mu_{i}^{*}([F,\partial F])=1$ for all $i=0,1,\ldots,n.$
    4. (d)
      Finally, to compute the values of $\delta h_{i}^{*}$ and $\delta\mu_{i}^{*}$ over $D_{i},\partial D_{i}]\in H_{2}(N_{i},T_{i})$, we note that $\partial[D_{i}]=-(a_{i}[h_{i}]+b_{i}[\mu_{i}])$ by our orientation convention in order to keep $b_{i}>0$ (see §3.1). Then it follows that for $i=0,1,\ldots,n$
      $\delta h_{i}^{*}([D_{i},\partial D_{i}])=h_{i}^{*}(-a_{i}[h_{i}]-b_{i}[\mu_{i}])=-a_{i}$
      and
      $\delta\mu_{i}^{*}([D_{i},\partial D_{i}])=\mu_{i}^{*}(-a_{i}[h_{i}]-b_{i}[\mu_{i}])=-b_{i}.$
In summary, we have
$c_{\sigma}]=\chi(F)F^{*}+\sum_{i=0}^{n}D_{i}^{*}$
The image of $\delta$ is generated by
$\delta\mu_{i}^{*}=F^{*}-\sum_{i=0}^{n}b_{i}D_{i}^{*},\quad\text{ and }\quad\delta h_{i}^{*}=\begin{cases}-c_{0}^{*}\otimes S^{*}-a_{0}D_{0}^{*},&\text{for }i=0,\\[6.0pt] c_{i-1}^{*}\otimes S^{*}-c_{i}^{*}\otimes S^{*}-a_{i}D_{i}^{*},&\text{for }0<i<n,\\[6.0pt] c_{n-1}^{*}\otimes S^{*}-a_{n}D_{n}^{*},&\text{for }i=n.\end{cases}$
if $n>0$; and when $n=0$, the image of $\delta$ is generated by
$\delta\mu_{0}^{*}=F^{*}-b_{0}D_{0}^{*}=F^{*}-D_{0}^{*}\quad\text{and}\quad\delta h_{0}^{*}=-a_{0}D_{0}^{*}=-bD_{0}^{*}.$
The Euler class $e(\nu_{M})=i^{*}([c_{\sigma}])\in H^{2}(M)$ vanishes if and only if $c_{\sigma}]\in\mathrm{Im}(\delta)$. That is, there exist integers $m_{i},n_{i}$, for $i=0,1,\ldots,n$, such that
$\delta\Bigl(\sum_{i}-m_{i}h_{i}^{*}+\sum_{i}n_{i}\mu_{i}^{*}\Bigr)=[c_{\sigma}].$
  1. (1)
    When $n>0$, for the coefficients of $c_{j}^{*}\otimes S^{*}$ to be zero, the $m_{i}$ must be identical. Hence it follows that $c_{\sigma}]\in\mathrm{Im}(\delta)$ if and only if there exist $m,n_{i}\in\mathbb{Z}$ for $i=0,\ldots,n$ such that
    $\begin{cases}ma_{i}-n_{i}b_{i}=1,\mbox{ for }i=0,1,\ldots,n\\[6.0pt] \sum_{i=0}^{n}n_{i}=\chi(F)=\chi(|\mathcal{B}|)-(n+1).\end{cases}$
    Recall that $\tfrac{a_{0}}{b_{0}}=b$. Hence the equation $ma_{0}-n_{0}b_{0}=1$ simplifies to $mb-n_{0}=1$, and therefore, the condition above is equivalent to the existence of $m,n_{i}\in\mathbb{Z}$ for $i=1,\ldots,n$ such that
    $\begin{cases}ma_{i}-n_{i}b_{i}=1,\text{ for }i=1,\ldots,n\\[6.0pt] mb-1+\displaystyle\sum_{i=1}^{n}n_{i}=\chi(F)=\chi(|\mathcal{B}|)-(n+1)\end{cases}$
    For $i\geq 1$, we have
    $n_{i}=\frac{ma_{i}-1}{b_{i}}=\frac{ma_{i}}{b_{i}}-\frac{1}{b_{i}}$
    Hence, the second condition is equivalent to
    $m(b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}})=\chi(|\mathcal{B}|)-n+\sum_{i=1}^{n}\frac{1}{b_{i}}$
    Note that
    $e(M)=b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}\quad\text{and}\quad\chi(\mathcal{B})=\chi(|\mathcal{B}|)-n+\sum_{i=1}^{n}\frac{1}{b_{i}}$
    This completes the proof of the theorem when $n>0$.
  2. (2)
    If $n=0$ (and hence there is no singular fibre), the condition above reduces to the existence of $m\in\mathbb{Z}$ such that
    $mb=\chi(\mathcal{B})$
    From the discussion above, one can easily verify that this is a necessary and sufficient condition for the Euler class to vanish in this special case.
This proves the theorem. ∎
4. Discrete representations of Fuchsian groups and the normal bundle of a Seifert fibration
In this section we provide alternative proofs of Theorem 1.11 and Corollary 1.12 from the perspective of discrete faithful representations of Fuchsian groups. We normalise our notation for oriented Seifert manifolds $M=M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ by further requiring that $\tfrac{a_{i}}{b_{i}}\in(0,1)$ for each $i=1,\ldots,n$. See Remark 3.1.
Figure 1. The generators $\bar{x}_{i}$ in the presentation of $\Gamma=\pi_{1}(\mathcal{B})$
The fundamental group of $M$ has a presentation
(4.0.1) $\hskip 28.45274pt\langle x_{1},x_{2},\ldots,x_{n},w_{1},w_{2},\ldots,w_{g},z_{1},z_{2},\ldots,z_{g},h\;|\;h\mbox{ central},x_{i}^{b_{i}}=h^{-a_{i}}\mbox{ for each }i,$
$\hskip 56.9055pt[w_{1},z_{1}][w_{2},z_{2}]\cdots[w_{g},z_{g}]x_{1}x_{2}\cdots x_{n}=h^{b}\rangle,$
and $\Gamma=\pi_{1}(\mathcal{B})$ has a presentation
(4.0.2) $\hskip 28.45274pt\langle\bar{x}_{1},\ldots,\bar{x}_{n},\bar{w}_{1},\ldots,\bar{w}_{g},\bar{z}_{1},\ldots,\bar{z}_{g}\;|\;\bar{x}_{i}^{b_{i}}=1,[\bar{w}_{1},\bar{z}_{1}]\cdots[\bar{w}_{g},\bar{z}_{g}]\bar{x}_{1}\cdots\bar{x}_{n}=1\rangle$
These two presentations are related via a well-known central extension:
(4.0.3) $E_{M}:1\to\mathbb{Z}\to\pi_{1}(M)\xrightarrow{\;\alpha\;}\Gamma\to 1,$
Here $h$ corresponds to the image in $\pi_{1}(M)$ of $1\in\mathbb{Z}$, $\alpha(x_{i})=\bar{x}_{i}$ for each $i$, $\alpha(w_{j})=\bar{w}_{j}$ for each $j$ and $\alpha(z_{k})=\bar{z}_{k}$ for each $k$.
4.1. The second cohomology group of $\Gamma$
Jankins and Neuman determined a canonical isomorphism between $H^{2}(\Gamma)$ and the abelian group having generators $A_{0},A_{1},\ldots,A_{n}$ and relations $A_{0}=b_{i}A_{i}$ for $1\leq i\leq n$ [16]. (The case that $g=0$ is discussed in [16] and the general case in [15].) Each element of $H^{2}(\Gamma)$ can be uniquely expressed in the form
$dA_{0}+c_{1}A_{1}+\cdots+c_{n}A_{n},$
where $d,c_{1},\ldots,c_{n}\in\mathbb{Z}$ and $1\leq c_{i}\leq b_{i}-1$ for each value of $i$.
The Euler class $e(E_{M})\in H^{2}(\Gamma)$ is the cohomology class which classifies the extension (4.0.3). In [16] it was shown to be
(4.1.1) $e(E_{M})=-(bA_{0}+a_{1}A_{1}+a_{2}A_{2}+\cdots+a_{n}A_{n})$
This class arises as the image of 1-1 under the transgression homomorphism $\delta$ from the $5$-term exact sequence in cohomology of $E_{M}$:
(4.1.2) $0\to H^{1}(\Gamma)\xrightarrow{}H^{1}(M)\to\mathbb{Z}\xrightarrow{\;\delta\;}H^{2}(\Gamma)\xrightarrow{\;\alpha^{*}\;}H^{2}(M)$
See [13, Theorem 4]. Thus,
(4.1.3) $\mbox{kernel}(\alpha^{*})=\langle e(E_{M})\rangle$
4.2. The normal bundle of the Seifert fibring of $M$ and discrete representations
Since $\mathcal{B}$ is hyperbolic, its universal cover can be identified with $\mathbb{H}^{2}$ and the associated group of deck transformations with a discrete subgroup of $\mbox{Isom}_{+}(\mathbb{H}^{2})=PSL_{2}(\mathbb{R})$. This yields a discrete faithful representation $\bar{\varphi}:\Gamma\to PSL_{2}(\mathbb{R})$.
There is a Seifert fibring $\widehat{M}\to\mathbb{H}^{2}$ lifting $M\to\mathcal{B}$ for which the following diagram commutes
$\widehat{M}=\mathbb{H}^{2}\times S^{1}$$\mathbb{H}^{2}$$M\mathcal{B}$
where $\widehat{M}$ as the total space of an $S^{1}$-bundle over the contractible $\mathbb{H}^{2}$ can be identified with $\mathbb{H}^{2}\times S^{1}$. Hence, we can identify the universal cover of $M$ with $\mathbb{H}^{2}\times\mathbb{R}$, where the vertical lines $\{x\}\times\mathbb{R}$ universally cover the Seifert fibres of $M$. Denote the normal bundle to these vertical lines by $\tilde{\nu}$.
The fundamental group of $M$ acts on $\mathbb{H}^{2}$ via the composition $\varphi=\bar{\varphi}\circ\alpha$, which gives rise to a diagonal action $\psi$ of $\pi_{1}(M)$ on $\widetilde{M}\times\mathbb{H}^{2}$:
$\psi(\gamma)(x,y)=(\gamma\cdot x,\varphi(\gamma)(M))$
This action is equivariant with respect to the projection $\widetilde{M}\times\mathbb{H}^{2}\to\widetilde{M}$, so taking quotients determines a locally-trivial $\mathbb{H}^{2}$-bundle
$M\times_{\psi}\mathbb{H}^{2}=(\widetilde{M}\times\mathbb{H}^{2})/\pi_{1}(M)\to M$
Fix a Riemannian metric on $M$ and pull it back to $\widetilde{M}$. Since the group of deck transformations of the cover $\widetilde{M}\to M$ acts as isometries of this metric, the associated exponential map $\mbox{exp}:T\widetilde{M}\to\widetilde{M}$ is $\pi_{1}(M)$-equivariant, as is the composition
$E(\tilde{\nu}_{M})\xrightarrow{\;\mbox{{\footnotesize exp}}\;}\widetilde{M}=\mathbb{H}^{2}\times\mathbb{R}\xrightarrow{\;\mbox{{\footnotesize projection}}\;}\mathbb{H}^{2}$
Now proceed as in the proof of [3, Lemma A.3], starting from its third paragraph, to show that the $\mathbb{H}^{2}$-bundle $M\times_{\psi}\mathbb{H}^{2}\to M$ is topologically equivalent to the normal bundle $\nu_{M}$ and conclude, as in [3, Proposition A4], that $e(\nu_{M})=e(\varphi)=e(\bar{\varphi}\circ\alpha)$. Thus
(4.2.1) $e(\nu_{M})=\alpha^{*}(e(\bar{\varphi}))$
It is well-known that $e(\bar{\varphi})$, respectively $e(\varphi)$, is the obstruction to lifting the representation $\bar{\varphi}:\Gamma\to PSL_{2}(\mathbb{R})$, respectively $\varphi:\pi_{1}(M)\to PSL_{2}(\mathbb{R})$, to $\widetilde{PSL}_{2}(\mathbb{R})$ [19, Lemma 2].
An expression for $e(\bar{\varphi})$ in terms of the Jankins-Neumann generators $A_{0},A_{1},\ldots,A_{n}$ of $H^{2}(\Gamma)$ can be determined following the method of §3 of [16]. To explain it, recall the central extension $1\to\mathbb{Z}\to\widetilde{PSL}_{2}(\mathbb{R})\to PSL_{2}(\mathbb{R})\to 1$. Let $h\in\widetilde{PSL}_{2}(\mathbb{R})$ correspond to $1\in\mathbb{Z}$.
By construction, the generator $\bar{x}_{i}$ of the presentation (4.0.2) of $\Gamma$ is mapped to a conjugate of a rotation of angle $\frac{2\pi}{b_{i}}$ in $PSL_{2}(\mathbb{R})\leq\mbox{Homeo}_{+}(\mathbb{S}^{1})$ by $\bar{\varphi}$, so lifts to a conjugate of translation by $\frac{1}{b_{i}}$ in $\widetilde{PSL}_{2}(\mathbb{R})\leq\mbox{Homeo}_{+}(\mathbb{R})$. Hence, if
$1\to\mathbb{Z}\to\widetilde{\Gamma}\to\Gamma\to 1$
is the pull-back central extension determined by $\bar{\varphi}$, then $\widetilde{\Gamma}$ has a presentation
$\langle\tilde{x}_{1},\ldots,\tilde{x}_{n},\tilde{w}_{1},\ldots,\tilde{w}_{g},\tilde{z}_{1},\ldots,\tilde{z}_{g}\;|\;\tilde{x}_{i}^{b_{i}}=h,[\tilde{w}_{1},\tilde{z}_{1}]\cdots[w_{g},\tilde{z}_{g}]\tilde{x}_{1}\cdots\tilde{x}_{n}=h^{-b}\rangle$
It is a classic result in the theory of Fuchsian groups that under our assumptions, $b=\chi(|\mathcal{B}|)-n$. As such, the discussion in §3 of [16] implies that
(4.2.2) $e(\bar{\varphi})=(n-\chi(|\mathcal{B}|))A_{0}-A_{1}-A_{2}-\cdots-A_{n}$
Proof of Theorem 1.11 when $\mathcal{B}$ is hyperbolic.
As $\nu_{M}$ is an oriented plane bundle, it is trivial if and only if $e(\nu_{M})=0$. Combining (4.2.1), (4.1.3), (4.1.1), and (4.2.2) shows that $e(\nu_{M})=0$ if and only if there is an integer $m$ such that
(4.2.3) $(\chi(|\mathcal{B}|)-n)A_{0}+A_{1}+\cdots+A_{n}=m(bA_{0}+a_{1}A_{1}+a_{2}A_{2}+\cdots+a_{n}A_{n})$
Consideration of the normal form for elements of $H^{2}(\Gamma)$ described in §4.1 shows that the existence of such an $m$ is equivalent to the existence of an integer $m$ for which the following two conditions hold
  1. (1)
    $ma_{i}\equiv 1\mbox{ (mod $b_{i}$) for }1\leq i\leq n$;
  2. (2)
    $mb+\sum_{i=1}^{n}\left\lfloor\frac{ma_{i}}{b_{i}}\right\rfloor=\chi(|\mathcal{B}|)-n$.
For each $i\in\{1,2,\ldots,n\}$, it follows from the congruence equation $ma_{i}\equiv 1\pmod{b_{i}}$ that
$\left\lfloor\frac{ma_{i}}{b_{i}}\right\rfloor=\frac{ma_{i}}{b_{i}}-\frac{1}{b_{i}}$
Hence, the second equation becomes
$m\left(b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}\right)=\chi(|\mathcal{B}|)-n+\sum_{i=1}^{n}\frac{1}{b_{i}}$
which again simplifies to
$me(M)=\chi(\mathcal{B})$
This completes the proof. ∎
Though we proved Corollary 1.12 as a consequence of Theorem 1.11 in §3.2, we provide a proof of the hyperbolic case of the corollary next which is independent of that theorem.
Proof of Corollary 1.12 when $\mathcal{B}$ is hyperbolic.
Let $l$ denote the least common multiple of the integers $b_{1},b_{2},\ldots,b_{n}$ and note that as $\gcd(\frac{l}{b_{1}},\frac{l}{b_{2}},\ldots,\frac{l}{b_{n}})=1$, we have an epimorphism
$\theta:H^{2}(\Gamma)\to\mathbb{Z},\;A_{0}\mapsto l,A_{i}\mapsto\frac{l}{b_{i}}\mbox{ for }1\leq i\leq n$
Then by (4.1.1),
(4.2.4) $\theta(e(E_{M}))=-l(b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}})=-le(M),$
where $e(M)$ is the Euler number of $M$ (see §3.1).
Next note that (4.2.2) implies that
(4.2.5) $\theta(e(\bar{\varphi}))=l(n-\chi(|\mathcal{B}|)-\sum_{i}\frac{1}{b_{i}})=-l\chi(\mathcal{B})$
Suppose that $e(\nu_{M})=0$. Then (4.2.1) and (4.1.3) imply that there is an integer $m$ such that $e(\bar{\varphi})=me(E_{M})\in H^{2}(\Gamma)$ and therefore $\theta(e(\bar{\varphi}))=m\theta(e(E_{M}))$. Since $\mathcal{B}$ is hyperbolic, $\chi(\mathcal{B})\neq 0$, and therefore (4.2.4) and (4.2.5) imply that $m\neq 0$ and $\chi(\mathcal{B})=me(M)$, which is what was to be proved. ∎
Proof of Theorem 1.14 when $\mathcal{B}$ is hyperbolic.
We saw in (4.2.1) that that $e(\nu_{M})=\alpha^{*}(e(\bar{\varphi}))=e(\bar{\varphi}\circ\alpha)=e(\varphi)$, where $\varphi=\bar{\varphi}\circ\alpha:\pi_{1}(M)\to PSL_{2}(\mathbb{R})$. Hence, if $e(\nu_{M})=0$, then $\varphi$ lifts to a homomorphism $\pi_{1}(M)\to\widetilde{PSL}_{2}(\mathbb{R})\leq\mbox{Homeo}_{+}(\mathbb{R})$ [19, Lemma 2]. It follows that $\pi_{1}(M)$ is left-orderable [8, Theorem 1.1] and therefore, $M$ admits a horizontal foliation [8, Theorem 1.3]. ∎
5. Seifert fibred examples
Recall that if $M$ is a Seifert fibred rational homology $3$-sphere, then the underlying surface of the base orbifold is either $S^{2}$ or $P^{2}$. However, in the latter case, $M$ does not admit a co-orientable taut foliation. (See [8] for instance.)
In this section we use Theorem 1.11 and Corollary 1.12 to analyse when an oriented Seifert manifold $M=M(0;b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$ has a trivial normal bundle and apply this to show that generically, if $M$ has a co-oriented taut foliation, it has none with zero Euler class.
We assume that $n\geq 3$. Then the base orbifold $\mathcal{B}$ of $M$ is spherical if $\chi(\mathcal{B})>0$, Euclidean if $\chi(\mathcal{B})=0$, and hyperbolic if $\chi(\mathcal{B})<0$. Orient $\mathcal{B}$ and endow $\nu_{M}$, the normal bundle to the Seifert fibres of $M$, with the induced orientation. The following lemma is key to our applications.
Lemma 5.1.
Suppose that $M$ is a Seifert fibred rational homology $3$-sphere. Then
$(1)$ Any co-orientable taut foliation on $M$ is horizontal. That is, it can be isotoped to be everywhere transverse to the Seifert fibres of $M$.
$(2)$ The tangent space to any co-orientable taut foliation on $M$ is isomorphic to $\nu_{M}$ as an unoriented bundle. In particular, the Euler class of any co-oriented taut foliation on $M$ agrees, up to sign, with that of $\nu_{M}$.
Proof.
That any co-oriented taut foliation on $M$ must be horizontal is [7, Corollary 7] in the case $n=3$ and [1, Proposition 6.1] in general. Thus (1) holds. Consequently, the tangent bundle of any such foliation is isomorphic to $\nu_{M}$ (as an unoriented bundle), which implies (2). ∎
Theorem 1.8 follows immediately from this lemma and Theorem 1.11.
Proof of Theorem 1.7.
Let $K$ be the right-handed trefoil knot, and let $\mu$ and $\lambda$ denote the standard meridian and longitude of $K$. There is a natural Seifert fibre structure on the exterior $X=S^{3}\setminus\nu(K)$, whose regular fibre has slope $6\mu+\lambda$ on $\partial X$. Hence, $p/q$-surgery on $K$ yields a small Seifert fibred space whenever $p/q\neq 6$.
It is known that $K(p/q)$ admits co-orientable taut foliations when $p/q<1$ [25, §3] and hence has a left-orderable fundamental group [4]. On the other hand, assuming that $p>0$ and $p,q$ are coprime, [14, Theorem 1.7] shows that the Euler class of the co-orientable taut foliation constructed by Roberts in [25] vanishes if and only if
$|q|\equiv 1\pmod{p}$.
This completes the proof. ∎
The following lemma is an immediate consequence of Theorem 1.11.
Lemma 5.2.
If $e(\nu_{M})=0$, then $d=\gcd(b_{i},b_{j})$ implies $a_{i}\equiv a_{j}\mbox{ {\rm(mod $d$)}}$.
Two sets of Seifert invariants $(b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$, $(b^{\prime},\frac{a_{1}^{\prime}}{b_{1}},\ldots,\frac{a_{n}^{\prime}}{b_{n}})$ are said to be equivalent, written
$(b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})\sim(b^{\prime},\frac{a_{1}^{\prime}}{b_{1}},\ldots,\frac{a_{n}^{\prime}}{b_{n}})$
if they differ by a finite sequence of moves each of which either multiplies the coordinates by 1-1 or is of the form described in Remark 3.1(1). Equivalent sets of Seifert invariants classify Seifert structures on a closed $3$-manifold up to homeomorphism. See Remark 3.1.
5.1. $\mathcal{B}$ is spherical
When $\mathcal{B}$ is spherical, $n=3$ and $(b_{1},b_{2},b_{3})$ is one of the Platonic triples $(2,2,b_{3})$, $(2,3,3)$, $(2,3,4)$, or $(2,3,5)$. In each case, $b_{1}=2$ and $b_{2}$ is either $2$ or $3$. The fundamental group of $M$ is finite in all cases and therefore $b_{1}b_{2}b_{3}e(M)=\pm|H_{1}(M)|\neq 0$ (i.e., $H_{1}(M)$ is finite). Further, $\chi(\mathcal{B})>0$.
By assumption, $a_{i}=1$ when $b_{i}=2$. If $b_{2}=3$ and $a_{2}=2$, then the Seifert invariants of $M$ are
$(b,\frac{1}{2},\frac{2}{3},\frac{a_{3}}{b_{3}})\sim(b+1,\frac{1}{2},\frac{-1}{3},\frac{a_{3}}{b_{3}})\sim(-b-1,-\frac{1}{2},\frac{1}{3},-\frac{a_{3}}{b_{3}})\sim(-b-3,\frac{1}{2},\frac{1}{3},\frac{b_{3}-a_{3}}{b_{3}})$
Thus we can assume that $M$ has Seifert invariants $(b,\frac{1}{2},\frac{1}{b_{2}},\frac{a}{b_{3}})$. The following table lists $\chi(\mathcal{B}),e(M)$ and $|H_{1}(M)|$ for these manifolds. The last column determines the Seifert invariants of $M$ when $e(\nu_{M})=0$ by applying Theorem 1.11 and Corollary 1.12.
Seifert invariants $\chi(\mathcal{B})$ $e(M)$ $\pm|H_{1}(M)|$ Seifert invariants when $e(\nu_{M})=0$
$(b,\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2})$ $\tfrac{1}{2}$ $\tfrac{2b+3}{2}$ $4(2b+3)$ $(-1,\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2})$ and $(-2,\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2})$
$(b,\tfrac{1}{2},\tfrac{1}{2},\tfrac{a}{b_{3}}),\,b_{3}\geq 3$ $\tfrac{1}{b_{3}}$ $\tfrac{b_{3}(b+1)+a}{b_{3}}$ $4(b_{3}(b+1)+a)$ $(-1,\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{b_{3}})$
$(b,\tfrac{1}{2},\tfrac{1}{3},\tfrac{1}{3})$ $\tfrac{1}{6}$ $\tfrac{6b+7}{6}$ $3(6b+7)$ $(-1,\tfrac{1}{2},\tfrac{1}{3},\tfrac{1}{3})$
$(b,\tfrac{1}{2},\tfrac{1}{3},\tfrac{a}{4})$ $\tfrac{1}{12}$ $\tfrac{12b+10+3a}{12}$ $2(12b+10+3a)$ $(-1,\tfrac{1}{2},\tfrac{1}{3},\tfrac{1}{4})$
$(b,\tfrac{1}{2},\tfrac{1}{3},\tfrac{a}{5})$ $\tfrac{1}{30}$ $\tfrac{30b+6a+25}{30}$ $30b+6a+25$ $(-1,\tfrac{1}{2},\tfrac{1}{3},\tfrac{1}{5})$
It is interesting to note that the values of the Seifert invariants when $e(\nu_{M})=0$ correspond to those which realise the minimal values of $|H_{1}(M)|$ in the given families.
5.2. $\mathcal{B}$ is Euclidean
Here, either $n=3$ and $(b_{1},b_{2},b_{3})$ is one of the triples $(2,3,6),(2,4,4)$, $(3,3,3)$ or $n=4$ and $(b_{1},b_{2},b_{3},b_{4})=(2,2,2,2)$. We can assume that $a_{i}=1$ in all cases. This is obvious when $b_{i}=2$ and using Lemma 5.2 and the arguments of §5.1 deals with the case that some $b_{i}>2$.
As $\mathcal{B}$ is Euclidean, its Euler characteristic is zero. Hence if $\nu_{M}$ is trivial, Corollary 1.12 implies that $e(M)=0$ and we can use this to determine the Seifert invariants of $M$ when $e(\nu_{M})=0$.
Seifert invariants $e(M)$ $\pm|H_{1}(M)|$ Seifert invariants when $e(\nu_{M})=0$
$(b,\tfrac{1}{2},\tfrac{1}{3},\tfrac{1}{6})$ $b+1$ $36(b+1)$ $(-1,\tfrac{1}{2},\tfrac{1}{3},\tfrac{1}{6})$
$(b,\tfrac{1}{2},\tfrac{1}{4},\tfrac{1}{4})$ $b+1$ $32(b+1)$ $(-1,\tfrac{1}{2},\tfrac{1}{4},\tfrac{1}{4})$
$(b,\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3})$ $b+1$ $27(b+1)$ $(-1,\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3})$
$(b,\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2})$ $b+2$ $16(b+2)$ $(-2,\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2})$
As in the spherical case, those $M$ with $e(\nu_{M})=0$ are characterised by their first homology groups having minimal order (0 in this case). Their first Betti numbers are positive and therefore they admit horizontal fibrations (by tori) over the circle. This proves Theorem 1.14 when $\mathcal{B}$ is Euclidean.
5.3. $\mathcal{B}$ is hyperbolic
We assume in this section that the Seifert invariants of
$M=M(0;b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$
have been normalised so that $0<a_{i}<b_{i}$ for $1\leq i\leq n$.
The results of [17] and [20] show that $M$ admits a horizontal foliation if and only if one of the following conditions holds:
  1. (1)
    $2-n\leq b\leq-2$;
  2. (2)
    b=1b=-1 and there are coprime integers $0<c<d$ such that for some permutation $(\frac{c_{1}}{d},\frac{c_{2}}{d},\ldots,\frac{c_{n}}{d})$ of $(\frac{c}{d},\frac{d-c}{d},\frac{1}{d},\ldots,\frac{1}{d})$ we have $\frac{a_{i}}{b_{i}}<\frac{c_{i}}{d}$ for each $i$.
  3. (3)
    b=1nb=1-n and there are coprime integers $0<c<d$ such that for some permutation $(\frac{c_{1}}{d},\frac{c_{2}}{d},\ldots,\frac{c_{n}}{d})$ of $(\frac{c}{d},\frac{d-c}{d},\frac{1}{d},\ldots,\frac{1}{d})$ we have $\frac{b_{i}-a_{i}}{b_{i}}<\frac{c_{i}}{d}$ for each $i$.
The following specific examples illustrate that there are many Seifert manifolds $M$ that have co-orientable taut foliations with non-zero Euler class.
Examples 5.3.
Using Lemma 5.2, if some $b_{i},b_{j}$ have a common factor greater than $2$, one can construct many examples $M$ with $e(\nu_{M})\neq 0$. On the other hand using the conditions (1), (2), (3) immediately above, one can arrange for many of these to have horizontal foliations.
In (1), (2), (3), and (4) below we give some other classes of examples of Seifert fibred manifolds which admit co-orientable taut foliations but none with vanishing Euler class. The proofs are all applications of Theorem 1.1; we omit the details.
(1) By Lemma 5.2, if $b_{1}=b_{2}=\cdots=b_{n}$ and $e(\nu_{M})=0$, then $a_{1}=a_{2}=\cdots=a_{n}$. The first set of examples deals with a special case of this situation.
For $n\geq 4$, consider the manifold $M=M(0;b,\frac{1}{c},...,\frac{1}{c})$. Then
  • (a)
    $M$ has a horizontal foliation if and only if either $2-n\leq b\leq-2$, or b=1b=-1 and $c\geq 3$.
  • (b)
    If $c\geq 2n-2$, then $e(\nu_{M})=0$ if and only if b=2nb=2-n.
The bound on $c$ in part (b) is optimal: if c=2n3c=2n-3 and b=1b=-1, then $e(\nu_{M})=0$.
(2) For $n\geq 3$, let $M=M(0;2-n,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$, where $\sum_{i}\frac{a_{i}}{b_{i}}\leq\frac{n-2}{2}$. Then $M$ has a horizontal foliation, but $e(\nu_{M})=0$ if and only if $a_{i}=1$ for $1\leq i\leq n$.
(3) For $n\geq 4$, consider the manifold $M=M(0;b,\frac{1}{b_{1}},\ldots,\frac{1}{b_{n}})$, where the $b_{i}$ are pairwise coprime and $\geq n$. Then
  • (a)
    $M$ has a horizontal foliation if and only if $2-n\leq b\leq-1$.
  • (b)
    $e(\nu_{M})=0$ if and only if b=2nb=2-n.
(4) For $n\geq 4$, let $M=M(0;b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$, where $\sum_{i}\frac{a_{i}}{b_{i}}\leq\frac{n}{6}$. Then
  • (a)
    $M$ has a horizontal foliation for $2-n\leq b\leq-2$, and for b=1b=-1 if $n=4,5$, or $6$.
  • (b)
    $e(\nu_{M})=0$ for at most one $b$ with $2-n\leq b\leq\frac{2-n}{2}$.